### v2

# Multi-Agent Reinforcement Learning from Human Feedback: Data Coverage and Algorithmic Techniques

Abstract
We initiate the study of Multi-Agent Reinforcement Learning from Human Feedback (MARLHF), exploring both theoretical foundations and empirical validations. We define the task as identifying Nash equilibrium from a preference-only offline dataset in general-sum games, a problem marked by the challenge of sparse feedback signals. Our theory establishes the upper complexity bounds for Nash Equilibrium in effective MARLHF, demonstrating that single-policy coverage is inadequate and highlighting the importance of unilateral dataset coverage. These theoretical insights are verified through comprehensive experiments. To enhance the practical performance, we further introduce two algorithmic techniques. (1) We propose a Mean Squared Error (MSE) regularization along the time axis to achieve a more uniform reward distribution and improve reward learning outcomes. (2) We utilize imitation learning to approximate the reference policy, ensuring stability and effectiveness in training. Our findings underscore the multifaceted approach required for MARLHF, paving the way for effective preference-based multi-agent systems.
Keywords multi-agent reinforcement learning  $\cdot$ reinforcement learning from human feedback  $\cdot$ dataset coverage
1 Introduction
Large language models (LLMs) have achieved significant progress in natural language interaction, knowledge acquisition, instruction following, planning and reasoning, which has been recognized as the sparks for AGI (Bubeck et al., 2023). The evolution of LLMs fosters the field of agent systems, wherein LLMs act as the central intelligence (Xi et al., 2023). In these systems, multiple LLMs can interact with each other as well as with external tools. For instance, MetaGPT assigns LLM agents various roles, akin to those in a technology company, enabling them to cooperate on complex software engineering tasks (Hong et al., 2023).
Despite some empirical successes in agent systems utilizing closed-source LLMs, finetuning these systems and aligning them with human preferences remains a challenge. Reinforcement learning from human feedback (RLHF) has played an important role in aligning LLMs with human preferences (Christiano et al., 2017; Ziegler et al., 2019). However, unexpected behavior can arise when multiple LLMs interact with each other. In addition, reward design has been a hard problem in multi-agent reinforcement learning (Devlin et al., 2011). Thus, it is crucial to further align the multi-agent system with human feedback.
We address this problem through both theoretical analysis and empirical experiments. Theoretically, we characterize the dataset coverage condition for MARLHF that enables learning the Nash equilibrium, which serves as a favorable policy for each player. Empirically, we validate our theoretical insights through comprehensive experiments utilizing the proposed algorithmic techniques.
1.1 Contributions and Technical Novelties
1. Necessary and Sufficient Dataset Coverage Condition for MARLHF.
In single-agent RLHF, (Zhu et al., 2023) demonstrated that single policy coverage is sufficient for learning the optimal policy. However, we prove that this condition no longer holds for MARLHF by providing a counterexample. Instead, we introduce an algorithm that operates under unilateral coverage, a condition derived from offline MARL (Cui and Du, 2022a; Zhong et al., 2022). Specifically, this condition requires the dataset to cover all unilateral deviations from a Nash equilibrium policy. For further details, see Section 4.
Figure 1: The overall pipeline of offline MARLHF. $\mathcal{D}$ is the preference dataset where $\tau$ is a one trajectory and $\mathbf{y}_{i}\in\{1,-1\}^{m}$ indicates which trajectory is preferred by each agent. $r_{\phi}$ is the learned reward. $\pi_{ref}$ is the learned reference policy using imitation learning.
2. Algorithmic Techniques for Practical Performance.
As a foundational exploration into MARLHF research, we focus on employing the simplest learning framework, incorporating only the essential techniques necessary to ensure the approach’s feasibility. The framework consists of three key components: 1) leveraging the preference dataset to learn a reward function, 2) mitigating extrapolation errors using a reference policy, and 3) determining the final policy. Figure 1 provides an overview of the process.
However, additional algorithmic techniques are needed in order to find a strong policy even if we have a dataset with good coverage according to the theory.
  • Reward regularization. We observed that the reward learned through standard Maximum Likelihood Estimation (MLE) is sparse and spiky, making it difficult for standard RL algorithms to utilize effectively (cf. Figure 2 (b2)). To address this, we introduce an additional Mean Squared Error (MSE) loss between the predictions of adjacent time steps as a form of regularization. This regularization helps to prevent the model from accumulating reward signals solely at the final time step or relying on reward-irrelevant observation patterns, which could otherwise result in the complete failure in producing meaningful predictions.
  • Imitation learning policy as a reference. We choose to follow the standard RLHF work (Wang et al., 2023b), where the extrapolation problem was alleviated by adding an extra KL reward to the final optimization objective. Since we do not have access to the actual reference model, we implement imitation learning across the entire dataset to approximate it. The final policy is trained using a DQN-based Value Decomposition Network (VDN) (Mnih et al., 2013; Sunehag et al., 2017). Our ablation study demonstrates the effectiveness of selecting an appropriate KL reward coefficient (see Table 3).
3. Experiment Results.
Our experiments, following the pipeline described above, confirm the theoretical necessity of unilateral coverage. We conducted comprehensive ablation studies on three cooperative Multi-Agent Particle Environment (MPE) scenarios (Mordatch and Abbeel, 2017): Spread-v3, Tag-v3, and Reference-v3. These studies focused on the hyperparameter selection for the reward regularization coefficient, KL reward coefficient, and dataset diversity. The empirical results (Table 2) demonstrate that: 1) simply adding trivial trajectories to expert demonstrations can enhance performance, 2) unilateral datasets are advantageous, and 3) dataset diversity contributes to lower variance.
Our ablation experiments underscore the effectiveness of the proposed algorithmic techniques. Additionally, we tested the sensitivity of the hyperparameters used in MSE and KL regularization, and we introduce a principled standardization technique that can efficiently tune hyperparameters across all environments and datasets.
2 Related Works
Reinforcement Learning from Human Feedback (RLHF).
RLHF (or preference-based RL) plays a pivotal role in alignment with various tasks such as video games (Warnell et al., 2018; Brown et al., 2019), robotics (Jain et al., 2013; Kupcsik et al., 2016; Christiano et al., 2023; Shin et al., 2023), image augmentation (Metcalf et al., 2024), and large language models (Ziegler et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022; Stiennon et al., 2022; Bai et al., 2022; Glaese et al., 2022; Ganguli et al., 2022; Ouyang et al., 2022). Additionally, a body of work focuses on the reward models behind preference data (Sadigh et al., 2017; Bıyık and Sadigh, 2018; Gao et al., 2022; Hejna and Sadigh, 2023). Recently, direct preference optimization (DPO, Rafailov et al. (2023)) and its variants (Azar et al., 2023; Rafailov et al., 2024) approach RLHF without directly handling the reward model. Theoretical studies have also explored guarantees, such as sample complexity and regret, and the limitations of certain RLHF algorithms (Novoseller et al., 2020; Xu et al., 2020; Pacchiano et al., 2023; Chen et al., 2022; Razin et al., 2023; Zhu et al., 2024a; Wang et al., 2023b; Xiong et al., 2024; Zhu et al., 2024b).
Offline Reinforcement Learning.
Offline RL (Lange et al., 2012; Levine et al., 2020) has achieved success in a wide range of real-world applications, including robotics (Pinto and Gupta, 2015; Levine et al., 2016; Chebotar et al., 2021; Kumar et al., 2023), healthcare (Raghu et al., 2017; Wang et al., 2018), and autonomous driving (Shi et al., 2021; Lee et al., 2024). Key algorithms such as Behavior Cloning, BRAC (Wu et al., 2019), BEAR (Kumar et al., 2019), and CQL (Kumar et al., 2020; Lyu et al., 2024) have driven these successes. Theoretical research on offline RL has primarily focused on sample complexity under various dataset coverage assumptions Le et al. (2019); Chen and Jiang (2019); Yin et al. (2020); Rashidinejad et al. (2023); Yin et al. (2021, 2022); Shi et al. (2022); Nguyen-Tang et al. (2022); Xie et al. (2022); Xiong et al. (2023); Li et al. (2024).
Multi-Agent Reinforcement Learning (MARL).
Many real-world scenarios are naturally modeled as multi-agent environments, whether cooperative or competitive. As a result, MARL has gained popularity in video games (Tian et al., 2017; Vinyals et al., 2017; Silver et al., 2017; Vinyals et al., 2019), network design (Shamsoshoara et al., 2018; Kaur and Kumar, 2020), energy sharing (Prasad and Dusparic, 2018), and autonomous driving (Palanisamy, 2019; Yu et al., 2020; Zhou et al., 2022). Prominent algorithms in MARL include IQL (Tan, 2003), MADDPG (Lowe et al., 2020), COMA (Foerster et al., 2017), MAPPO (Yu et al., 2022), VDN (Sunehag et al., 2017), and QMIX (Rashid et al., 2018).
Offline MARL.
Offline MARL is a practical solution for handling sophisticated multi-agent environments. Empirically, to address issues related to out-of-distribution actions and complex reward functions, previous works have developed algorithms such as MABCQ (Jiang and Lu, 2023), ICQ-MA (Yang et al., 2021), OMAR (Pan et al., 2022), and OMIGA (Wang et al., 2023a), which incorporate regularization or constraints on these actions and functions. MOMA-PPO (Barde et al., 2024) is a model-based approach to offline MARL that generates synthetic interaction data from offline datasets. Tseng et al. (2022) combines knowledge distillation with multi-agent decision transformers (Meng et al., 2022) for offline MARL. Theoretical understanding of offline MARL, particularly in the context of Markov games, has been advanced by works that provide sample complexity guarantees for learning equilibria Sidford et al. (2019); Cui and Yang (2020); Zhang et al. (2023a, 2020); Abe and Kaneko (2020); Cui and Du (2022a, b); Zhang et al. (2023b); Blanchet et al. (2023); Shi et al. (2023).
3 Preliminaries
General-sum Markov Games. We consider an episodic time-inhomogeneous general-sum Markov game $\mathcal{M}$, consisting of $m$ players, a shared state space $\mathcal{S}$, an individual action space $\mathcal{A}_{i}$ for each player $i\in[m]$ and a joint action space $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\mathcal{A}_{m}$. The game has a time horizon $H$, an initial state $s_{1}$, state transition probabilities $\mathbb{P}=(\mathbb{P}_{1},\mathbb{P}_{2},\cdots,\mathbb{P}_{H})$ with $\mathbb{P}_{h}:\mathcal{S}\mathcal{A}\rightarrow\Delta(\mathcal{S})$, and rewards $R={R_{h}(\cdot\mid s_{h},a_{h})}_{h=1}^{H}$ where $R_{h,i}\in[0,1]$ represents the random reward for player $i$ at step $h$. At each step $h\in[H]$, all players observe current state $s_{h}$ and simultaneously choose their actions $\mathbf{a}_{h}=(a_{h,1},a_{h,2},\cdots,a_{h,m})$. The next state $s_{h+1}$ is then sampled from $\mathbb{P}_{h}(\cdot\mid s_{h},\mathbf{a}_{h})$, and the reward $r_{h,i}$ for player $i$ is sampled from $R_{h,i}(\cdot\mid s_{h},\mathbf{a}_{h})$. The game terminates at step $H+1$, with each player aiming to maximize the total collected rewards.
We use $\pi=(\pi_{1},\pi_{2},\cdots,\pi_{m})$ to denote a joint policy, where the individual policy for player $i$ is represented as $\pi_{i}=(\pi_{1,i},\pi_{2,i},\cdots,\pi_{H,i})$, with each $\pi_{h,i}:S\rightarrow\Delta(A_{i})$ defined as the Markov policy for player $i$ at step $h$. The state value function and state-action value function for each player $i\in[m]$ are defined as
$V_{h,i}^{\pi}(s_{h}):=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H}r_{t,i}(s_{t},\mathbf{a}_{t})\mid s_{h}\right],\ Q_{h,i}^{\pi}(s_{h}):=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H}r_{t,i}(s_{t},\mathbf{a}_{t})\mid s_{h},\mathbf{a}_{h}\right],$
where $\mathbb{E}_{\pi}=\mathbb{E}_{s_{1},\mathbf{a}_{1},\mathbf{r}_{1},\cdots,s_{H+1}\sim\pi,\mathcal{M}}$ denotes the expectation over the random trajectory generated by policy $\pi$. The best response value for player $i$ is defined as
$V_{h,i}^{\dagger,\pi_{-i}}(s_{h}):=\max_{\pi_{i}}V_{h,i}^{\pi_{i},\pi_{-i}}(s_{h}),$
which represents the maximal expected total return for player $i$ given that the other players follow policy $\pi_{-i}$.
A Nash equilibrium is a policy configuration where no player has an incentive to change their policy unilaterally. Formally, we measure how closely a policy approximates a Nash equilibrium using the Nash-Gap:
[…]
Offline Multi-agent Reinforcement Learning with Preference Feedback. In offline MARL with Preference Feedback, the algorithm has access to a pre-collected preference dataset generated by an unknown behavior policy interacting with an underlying Markov game. We consider two sampled trajectories, $\tau=(s_{1},\mathbf{a}_{1},s_{2},\mathbf{a}_{2},\cdots,s_{H+1})$ and $\tau^{\prime}=(s^{\prime}_{1},\mathbf{a}^{\prime}_{1},s^{\prime}_{2},\mathbf{a}^{\prime}_{2},\cdots,s^{\prime}_{H+1})$, drawn from distribution $\mathbb{P}(s_{1},\mathbf{a}_{1},s_{2},\cdots,s_{H+1})=\Pi_{h}\pi^{b}(\mathbf{a}_{h}\mid s_{h})\mathbb{P}(s_{h+1}\mid s_{h},\mathbf{a}_{h})$ induced by the behavior policy $\pi^{b}$. In MARLHF, the reward signal is not revealed in the dataset. Instead, each player can observe a binary signal $y_{i}$ from a Bernoulli distribution following the Bradley-Terry-Luce model (Bradley and Terry, 1952):
$\mathbb{P}(y_{i}=1\mid\tau,\tau^{\prime})=\frac{\exp(\sum_{h=1}^{H}r_{i}(s_{h},\mathbf{a}_{h}))}{\exp(\sum_{h=1}^{H}r_{i}(s_{h},\mathbf{a}_{h}))+\exp(\sum_{h=1}^{H}r_{i}(s^{\prime}_{h},\mathbf{a}^{\prime}_{h}))},\forall i\in[m].$
We make the standard linear Markov game assumption (Zhong et al., 2022):
Assumption 1.
$\mathcal{M}$ is a linear Markov game with a feature map $\psi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ if we have
$\mathbb{P}_{h}(s_{h+1}\mid s_{h},\mathbf{a}_{h})=\left\langle\psi(s_{h},\mathbf{a}_{h}),\mu_{h}(s_{h+1})\right\rangle,\forall(s_{h},\mathbf{a}_{h},s_{h+1},h)\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}\times[H],$
$r_{i}(s_{h},\mathbf{a}_{h})=\left\langle\psi(s_{h}),\theta_{h,i}\right\rangle,\forall(s_{h},\mathbf{a}_{h},h,i)\in\mathcal{S}\times\mathcal{A}\times[H]\times[m],$
where $\mu_{h}$ and $\theta_{h,i}$ are unknown parameters. Without loss of generality, we assume $\left\|\psi(s,\mathbf{a})\right\|\leq 1$ for all $(s,\mathbf{a})\in\mathcal{S}\times\mathcal{A}$ and $\left\|\mu_{h}(s)\right\|\leq\sqrt{d},\left\|\theta\right\|_{h}\leq\sqrt{d}$ for all $h\in[H]$.
The one-hot feature map is defined as $\overline{\psi}_{h}(s,\mathbf{a}):=[0,\cdots,0,\psi(s,\mathbf{a}),0,\cdots,0]\in\mathbb{R}^{Hd}$, where $\psi(s,\mathbf{a})$ is at position $(h-1)d+1$ to $hd$.
Value-Decomposition Network (VDN). In our experiments, we utilize VDN as an offline MARL algorithm for its effectiveness and simplicity. VDN (Sunehag et al., 2017) is a Q-learning style MARL architecture for cooperative games. It takes the idea of decomposing the team value function into agent-wise value functions, expressed as: $Q_{h}(s,\mathbf{a})=\sum_{i=1}^{n}Q_{h,i}(s,a_{i}).$ In our experiments, we applied Deep Q-Network (DQN)(Mnih et al., 2013) with VDN to learn the team Q function. We chose DQN to maintain the simplicity and controllability of the experimental pipeline, which facilitates a more accurate investigation of the impact of various techniques on the learning process.
4 Dataset Coverage Theory for MARLHF
In this section, we study the dataset coverage assumptions for offline MARLHF. For offline single-agent RLHF, (Zhu et al., 2023; Zhan et al., 2023) shows that single policy coverage is sufficient for learning the optimal policy. However, we prove that this assumption is insufficient in the multi-agent setting by constructing an counterexample. In addition, we prove that unilateral policy coverage is adequate for learning the Nash equilibrium.
4.1 Policy Coverages
We quantify the information contained in the dataset using covariance matrices, as the rewards and transition kernels are parameterized by a linear model. With a slight abuse of the notation, for trajectory $\tau=(s_{1},\mathbf{a}_{1},s_{2},\mathbf{a}_{2},\cdots,s_{H+1})$, we use $\psi(\tau):=[\psi(s_{1},\mathbf{a}_{1}),\psi(s_{2},\mathbf{a}_{2}),\cdots,\psi(s_{H},\mathbf{a}_{H})]$ to denote the concatenated trajectory feature. The reward coverage is measured by the preference covariance matrix:
$\Sigma_{\mathcal{D}}^{r}=\lambda I+\sum_{(\tau,\tau^{\prime})\in\mathcal{D}}(\psi(\tau)-\psi(\tau^{\prime}))(\psi(\tau)-\psi(\tau^{\prime}))^{\top},$
where $\psi(\tau)-\psi(\tau^{\prime})$ is derived from the preference model. Similarly, the transition coverage is measured by the covariance matrix:
$\Sigma_{\mathcal{D},h}^{\mathbb{P}}=\lambda I+\sum_{(\tau,\tau^{\prime})\in\mathcal{D}}\left[\psi(s_{h},\mathbf{a}_{h})\psi(s_{h},\mathbf{a}_{h})^{\top}+\psi(s^{\prime}_{h},\mathbf{a}^{\prime}_{h})\psi(s^{\prime}_{h},\mathbf{a}^{\prime}_{h})^{\top}\right].$
For a given state and action pair $(s_{h},\mathbf{a}_{h})$, the term $\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}]^{-1}}$ measures the uncertainty in reward estimation and $\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}]^{-1}}$ measures the uncertainty in transition estimation. As a result, the overall uncertainty of a given policy $\pi$ with dataset $\mathcal{D}$ is measured by
$U_{\mathcal{D}}(\pi):=\mathbb{E}_{\pi}\left[\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},a_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}]^{-1}}+\sum_{h=1}^{H}\left\|\psi(s_{h},a_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}]^{-1}}\right].$
Definition 1.
For a Nash equilibrium $\pi^{*}$, different policy coverages are measured by the following quantities:
  • Single policy coverage: $U_{\mathcal{D}}(\pi^{*})$.
  • Unilateral policy coverage: $\max_{i,\pi_{i}}U_{\mathcal{D}}(\pi_{i},\pi_{-i}^{*})$.
  • Uniform policy coverage: $\max_{\pi}U_{\mathcal{D}}(\pi)$.
Intuitively, small $U_{\mathcal{D}}(\pi^{*})$ indicates that the dataset contains adequate information about $\pi^{*}$. A small $\max_{i,\pi_{i}}U_{\mathcal{D}}(\pi_{i},\pi_{-i}^{*})$ implies that the dataset covers all of the unilateral deviations of $\pi^{*}$, and small $\max_{\pi}U_{\mathcal{D}}(\pi^{*})$ suggests that the dataset covers all possible policies.
4.2 Single Policy Coverage is Insufficient
Our objective is to learn a Nash equilibrium policy from the dataset, which necessitates that the dataset sufficiently covers the Nash equilibrium. In the single-agent scenario, if the dataset covers the optimal policy, pessimism-based algorithms can be employed to recover the optimal policy. However, previous work (Cui and Du, 2022a; Zhong et al., 2022) has demonstrated that single policy coverage is insufficient for offline MARL. We extend this result to the context of offline MARL with preference feedback, as follows:
Theorem 1.
(Informal) If the dataset only has coverage on the Nash equilibrium policy (i.e. small $U_{\mathcal{D}}(\pi^{*})$), it is not sufficient for learning an approximate Nash equilibrium policy.
The proof is derived by a reduction from standard offline MARL to MARLHF. Suppose that MARLHF with single policy coverage suffices, we could construct an algorithm for standard offline MARL, which leads to a contradiction. The formal statement and the detailed proof are deferred to Appendix A.1.
4.3 Unilateral Policy Coverage is Sufficient
While single policy coverage is too weak to learn a Nash equilibrium, uniform policy coverage, though sufficient, is often too strong and impractical for many scenarios. Instead, we focus on unilateral policy coverage, which offers a middle ground between single policy coverage and uniform policy coverage.
Theorem 2.
(Informal) If the dataset has unilateral coverage on the Nash equilibrium policy, there exists an algorithm that can output an approximate Nash equilibrium policy.
The detailed proof is deferred to Appendix A.2. We leverage a variant of Strategy-wise Bonus and Surrogate Minimization (SBSM) algorithm in (Cui and Du, 2022b) with modified policy evaluation and policy optimization subroutines. Intuitively, the algorithm identifies a policy that minimizes a pessimistic estimate of the Nash gap. As a result, if the dataset has unilateral coverage, the output policywill have a small Nash gap and serves as a good approximation of the Nash equilibrium.
5 Algorithmic Techniques for Practical Performance
In Section 4, we provided a theoretical characterization of the dataset requirements for MARLHF. However, the algorithm used in Theorem 2 is not computationally efficient. In this section, we propose a practical algorithm for MARLHF and validate our theoretical findings through experiments.
5.1 High-level Methodology
Our MARLHF pipeline consists of two phases: In the first step, we train a reward prediction model $\phi$ and approximate the behavior policy $\pi^{b}$ using imitation learning; in the second step, we then apply an MARL algorithm to maximize a combination of the KL-divergence-based reward and standardized predicted reward $r_{\phi}$, ultimately deriving the final policy $\pi_{\mathbf{w}}$.
Step 1: Reward Training and Reference Approximation. Given the preference signals of trajectories, we use neural networks to predict step-wise rewards $r_{\phi}(s_{h},a_{h})$ for each agent, minimizing the loss defined in (1). The objective is to map $(s,a_{i})$-pairs to reward values such that the team returns align with the preference signals. At the same time, in order to utilize KL reward $-\log\frac{\pi_{\mathbf{w}}(s,a)}{\pi^{b}(s,a)}$ to cope with the extrapolation error in offline learning, an imitation learner $\pi_{\text{ref}}$ is trained over the entire dataset to model the behavior policy $\pi^{b}$.
Step 2: Offline MARL. Although in this work, VDN is chosen as the MARL oracle, it should be noted that other MARL architectures are also applicable. With the reward model $r_{\phi}$ and the approximated reference policy $\pi_{\text{ref}}$ learned in the first step, we are now able to construct a virtual step-wise reward for each agent. The agents are then trained to maximize the target defined in (3).
Given this framework, additional techniques are required to build a strong practical algorithm, which we provide more details below.
5.2 Reward Regularization
Compared to step-wise reward signals, preference signals are $H$ times sparser, making them more challenging for a standard RL algorithm to utilize effectively. Concretely, this reward sparsity causes the naive optimization of the negative log-likelihood (NLL) loss to suffer from two key problems:
  1. 1.
    Sparse and spiky reward output. When calculating NLL losses, spreading the reward signal along the trajectories is equivalent to summing it at the last time step (Figure  2(a)). However, a sparse reward signal is harder for traditional RL methods to handdle due to the lack of continuous supervision. More uniformly distributed rewards across the entire trajectory generally leads to more efficient learning in standard RL algorithms.
  2. 2.
    Over-reliance on irrelevant features. The model may exploit redundant features as shortcuts to predict rewards. For instance, expert agents in cooperative games usually exhibit a fixed pattern of collaboration from the very beginning of the trajectory (such as specific actions or communication moves). The reward model might use these patterns to differentiate them from agents of other skill levels, thereby failing to capture the true reward-observation causal relationships.
To mitigate these problems, we introduce an extra Mean Squared Error (MSE) regularization along the time axis (Equation 1, 2). By limiting the sudden changes in reward predictions between adjacent time steps, this regularization discourages the reward model from concentrating its predictions on just a few time steps. While these issues can also be mitigated by using more diversified datasets and adding regularization to experts to eliminate reward-irrelevant action patterns, these approaches can be costly and sometimes impractical in real-world applications. In contrast, our MSE regularization is both easy to implement and has been empirically verified to be effective, creating more uniform reward distribution (Figure 2) and better performances.
$L_{\text{RM}}(\phi)=-\mathbb{E}_{\mathcal{D}}\left[\sum_{i=1}^{m}\log\sigma(y_{i}(r_{\phi,i}(\tau_{1})-r_{\phi,i}(\tau_{2})))\right]+\frac{\alpha}{\text{Var}_{\mathcal{D}}(r_{\phi})}L_{\text{MSE}}(\phi,\tau),$ (1)
where the regularization term $L_{\text{MSE}}$ is defined as:
$L_{\text{MSE}}(\phi,\tau)=\mathbb{E}_{\mathcal{D}}\left[\sum_{h=1}^{H-1}\|r_{\phi}(s_{h},\mathbf{a}_{h})-r_{\phi}(s_{h+1},\mathbf{a}_{h+1})\|_{2}^{2}\right].$ (2)
Here $\alpha$ is the regularization coefficient, which is set to be 1 in our experiments. The variance of $r_{\phi}$ is calculated over the training set to adaptively scale the regularization term. During training, $\text{Var}_{\mathcal{D}}(r_{\phi})$ is detached to prevent gradients from flowing through it. The effectiveness of this method is validated in the ablation study (cf. Section 6.3).
5.3 Imitation Learning Policy as Reference
There are various methods to mitigate the over-extrapolation errors in offline RL (Peng et al., 2019; Nair et al., 2021), including conservative loss over the Q-function (Kumar et al., 2020) and directly restricting the learned policy actions to those within within the dataset (Fujimoto et al., 2019). For simplicity and consistency with the former RLHF framework (Ouyang et al., 2022), we adopt the same per-step KL reward for enforcing restrictions between $\pi^{b}$ and $\pi_{\mathbf{w}}$. In many scenarios, a direct reference to the behavior policy is not accessible. As an alternative, imitation learning is used to estimate the reference policy, $\pi_{\text{ref}}\approx\pi^{b}$. To stabilize training, we standardize $r_{\phi}$ over $\mathcal{D}$ before combining it with the KL reward to make it comparable:
$\text{objective}(\mathbf{w})=\mathbb{E}_{\tau\sim\pi_{\mathbf{w}}}\left[\sum_{h=1}^{H}r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)-\beta\log\frac{\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})}{\pi_{\text{ref}}(s_{h},\mathbf{a}_{h})}\right],$ (3)
where $\beta$ is the KL reward coefficient, set to be (1, 1, 3) in Spread-v3, Reference-3 and Tag-v3 respectively, and the standardized reward $r_{\text{std}}$ is defined as:
$r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)=\sum_{i=1}^{m}\frac{r_{\phi}(s_{h},a_{h,i})-\mathbb{E}_{\mathcal{D}}({r}_{\phi})}{\sqrt{\text{Var}_{\mathcal{D}}(r_{\phi})}}.$ (4)
Notice that the explicit expression of $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ only exists in methods with an actor model, which leads to certain planning-based algorithms, such as Q-learning methods, being unable to directly use Equation 3. To enhance the compatibility of our pipeline, we substitute $1/|A|$ for $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ in such cases. Although this substitution formally removes the KL reward, in algorithms like Q-learning that output deterministic policies, the replaced reward is equivalent to encouraging the agent to match its deterministic choices with the actions most preferred by the reference policy. In other words, it seeks the closet element in the deterministic policy family to the reference policy, which is a degradation of the KL reward restricted in the domain of deterministic policy. The effectiveness of this method is validated in the ablation study (cf. Section 6.4).
6 Experiments
We design a series of experiments to validate our theories and methods in common general-sum games. Specifically, we first use VDN Sunehag et al. (2017) to train expert agents, and take intermediate checkpoints as rookie agents. Then, we use these agents to collect datasets and use the Broadley-Terry model over standardized returns to simulate human preference. Experiments are carried out to verify the efficiency of our approach with unilateral policy dataset coverage (in Theorem 2) while single policy coverage is insufficient (stated in Theorem 1). We also design ablation studies to showcase the importance of our methods, particularly focusing on reward structure regularization and imitation poclies as reference models.
6.1 Environments
Our experiments involved 3 Multi-Agent Particle Environments (MPE) implemented with JaxMARL codebase (Rutherford et al., 2023). Specifically, we utilized the Spread, Tag, and Reference scenarios, which covers three different levels of communication requirements. Spread-v3 contains a group of agents and target landmarks, where the objective is to cover as many landmarks as possible while avoiding collisions. Tag-v3 contains two opposing groups, where quicker "preys" need to escape from "predators". To ensure a fair comparison of different predator cooperation policies, we fixed a pretrained prey agent. Reference-v3 involves two agents and potential landmarks, where the agents need to find each one’s target landmark to receive a high reward. The target landmark of each agent is only known by the other agent at first, necessitating communication before they can move toward the correct landmark. A more detailed description of the tasks and their associated challenges is provided in Appendix B.2.
6.2 The Importance of Dataset Diversity
To study the influence of diversity of dataset, we manually designed 4 kinds of mixed joint behavior polices, and change their ratios to form different datasets.
  • Expert policy: $n$ expert agents. Trained with VDN till convergence.
  • Rookie policy: $n$ rookie agents. Trained with VDN with early stop.
  • Trivial policy: $n$ random agents. All actions are uniformly sampled from the action space.
  • Unilateral policy: $n-1$ expert agents and $1$ rookie agent of different proficiency level.
Table 1 presents the ratio of trajectories collected by the four different policies. The experiments are designed to hierarchically examine the roles of diversity (Diversified vs. Mix-Unilateral), unilateral coverage (Mix-Unilateral vs. Mix-Expert), and trivial comparison (Mix-Expert vs. Pure-Expert).
The ranking of diversity follows the order:
[…]
The experimental results are presented in Table 2. We can clearly observe that trivial comparisons (Mix-Expert vs. Pure-Expert) play a crucial role in learning an appropriate reward. A dataset containing only expert policies proves hard for the reward model to learn from the very beginning. In more challenging tasks, as reflected by higher MSE, the importance of unilateral coverage and diversity becomes more pronounced. Suboptimal cooperative agent actions assist the reward model in identifying the Nash equilibrium, thus highlighting key reward signals more effectively, as suggested by Theorem 2. Additionally, we found that in a diversified dataset, the variance across all training runs was lower compared to other datasets.
Beyond test returns, we also visualized the reward predictions and utilized the MSE between standardized reward predictions and ground truth as a measure of prediction quality. Although this MSE is not necessarily directly correlated with final performance—since accurate predictions do not need to match the shape of the ground truth—it can serve as a useful empirical metric.
Expert Unilateral Rookie Trivial
Diversified 1 1 1 1
Mix-Unilateral 2 1 0 1
Mix-Expert 3 0 0 1
Pure-Expert 4 0 0 0
Table 1: Final datasets mixed with various ratios. The overall dataset size is kept to 38400 trajectories. More details list in Appendix B.1.
Spread-v3 Tag-v3 Reference-v3 Spread-v3 Tag-v3 Reference-v3
Return Return Return MSE MSE MSE
Diversified -21.63 $\pm$ 0.51 28.94 $\pm$ 1.04 -19.59 $\pm$ 0.55 0.51 1.04 0.55
Mix-Unilateral -20.98 $\pm$ 0.56 32.04 $\pm$ 1.12 -20.60 $\pm$ 1.15 0.56 1.12 1.15
Mix-Expert -21.11 $\pm$ 1.16 33.01 $\pm$ 0.70 -20.38 $\pm$ 1.06 1.16 0.70 1.06
Pure-Expert -21.45 $\pm$ 1.01 40.14 $\pm$ 0.75 -28.19 $\pm$ 0.92 1.01 0.75 0.92
Table 2: Test returns and the mean squared error (MSE) between the standardized predicted rewards and the standardized ground truth rewards. Detailed descriptions of the datasets are provided in Section 6.2. In environments with smoother reward distributions, such as Spread-v3 and Reference-v3, dataset diversity primarily reduces variances. However, in more challenging environments, such as Tag-v3, dataset diversity plays a substantially more significant role.
6.3 Experiments for Reward Regularization
In Figure 2, we examined the effectiveness of our proposed reward regularization technique. Figure 2(a) demonstrates that without regularization, the learned rewards tend to be sparse and spiky compared to the ground truth rewards. The predicted rewards in (b1) and (b3) were learned using our reward regularization, while those in (b2) were learned without it ($\alpha=0$). Our results indicate that the reward regularization technique produces much smoother reward functions. We also quantitatively compare learned rewards with and without using reward regularization in finding the policy (cf. the first and the last column of Table 3).
We also observe that the rewards often exhibit temporal continuity, which can create greater discrepancies with the sparse, pulse-like ground truth. Notably, we found that adding stronger regularization does not necessarily lead to underfitting of the reward model; in some cases, it even helps the model converge to a lower training loss. Detailed parameters and experimental results are provided in the appendix (cf. Table 7). We attribute this to the role of regularization in preventing the model from overly relying on shortcuts.
(a) Mix-Expert $\alpha=0$ (spread-v3)
(b) reference-v3
(c) (b1) Pure-Expert    (b2) Diversified $\alpha=0$    (b3) Diversified
Figure 2: (a) Averaged reward predictions and ground truth of a trajectory sample on spread-v3. (b) Standardized reward predictions and ground truth of a trajectory sample in reference-v3. When trained with expert data only (b1), $\phi$ experiences a mode collapse, failing to give informative signals. Reward function trained without regularization (b2) shows spiky patterns and tends to accumulate predictions at certain time steps when trained with less diversified datasets as (a). Our method with diversified dataset (b3) gives predictions that approximate the ground truth well.
$\beta=0$ $\beta=0.1$ $\beta=1$ $\beta=10$ $\beta=100$ $\alpha=0$
Spread-v3 -22.56 $\pm$ 1.61 -22.03 $\pm$ 0.67 -20.82 $\pm$ 0.53 -20.46 $\pm$ 0.51 -20.35 $\pm$ 0.43 -22.21 $\pm$ 0.72
Tag-v3 4.11 $\pm$ 1.66 4.25 $\pm$ 0.53 10.96 $\pm$ 1.20 28.88 $\pm$ 1.02 29.53 $\pm$ 1.35 30.77 $\pm$ 0.57
Reference-v3 -19.69 $\pm$ 0.36 -19.37 $\pm$ 0.53 -18.89 $\pm$ 0.78 -18.33 $\pm$ 0.42 -18.54 $\pm$ 0.46 -21.86 $\pm$ 0.73
Table 3: Comparison of test return per agent with different parameters. Standard pipeline take KL reward coefficient $\beta=1$ for Spread-v3, Reference-v3 and $\beta=10$ for Tag-v3, and the MSE reward regularization coefficient $\alpha$ is set to the optimal value for fixed $\beta$. All the agents are trained on Diversified Dataset across 5 random seeds. This experiment shows the importance using reward regularization and reference policy.
6.4 Experiments for Imitation Learning
Intuitively, the choice of the KL coefficient $\beta$ depends on the quality of the reward model $\phi$ and the dataset. As shown in Table 3, simpler environments (e.g., Spread and Reference) require less KL restriction, while harder environments (e.g., Tag) need stronger restriction to avoid suffering from the extrapolation problem. A large $\beta$ can amplify the approximation error of imitation learning, potentially overshadowing the reward signal, whereas a small $\beta$ may fail to introduce enough pessimism to effectively mitigate extrapolation errors. In environments with relatively simple reward structures, such as Spread-v3 and Reference-v3, the standard choice of $\beta=1$ is near optimal. However, in more complex scenarios where the reward is harder to derive, increasing $\beta$ to 3 yields better performance.
7 Discussion
In this paper, we proposed dedicated algorithmic techniques for offline MARLHF and provided theoretical justification for the unilateral dataset coverage condition. We believe our work is a significant step towards systematically studying MARLHF and offers a foundational framework for future research in this area. The flexibility of our framework allows for application across a wide range of general games, and our empirical results validate the effectiveness of our proposed methods in various scenarios.
Looking ahead, there is significant potential to extend this work to more complex, real-world scenarios, particularly by integrating Large Language Models (LLMs) into multi-agent systems. Future research will focus on fine-tuning and aligning LLMs within MARLHF, addressing challenges such as increased complexity and the design of effective reward structures.
Appendix A Missing Proofs in Section 4
A.1 Single Policy Coverage is Insufficient
$a_{1}$ $a_{2}$
$b_{1}$ 0.5 1
$b_{2}$ 0 0.5
$a_{1}$ $a_{2}$
$b_{1}$ 0.5 0
$b_{2}$ 1 0.5
Table 4: Here we present two matrix games, $\mathcal{M}_{1}$ (left) and $\mathcal{M}_{2}$ (right). The row player aims to maximize their reward, while the column player aims to minimize it. The Nash Equilibrium in $\mathcal{M}_{1}$ is $(a_{1},b_{1})$, and in $\mathcal{M}_{2}$ it is $(a_{2},b_{2})$. Note that with a dataset covering only these two states, it is impossible to distinguish between these two games, and therefore, it is not possible to identify the exact Nash Equilibrium.
Theorem 3.
(Restatement of Theorem 1) For any algorithm and constant $C>0$, there exists a Markov game and a compliant dataset with $U_{\mathcal{D}}(\pi^{*})\leq C$ such that the output policy is at most an $0.5$-Nash equilibrium.
Proof.
We construct two linear Markov games with a shared compliant dataset such that no policy is a good approximate Nash equilibrium in both Markov games. Similar to [Cui and Du, 2022a], we consider Markov games with $H=1$, $m=2$, $\mathcal{A}_{1}=\{a_{1},a_{2}\}$ and $\mathcal{A}_{2}=\{b_{1},b_{2}\}$ with deterministic reward presented in Table 4.
The feature mapping for these two games is
$\psi(a_{1},b_{1})=e_{1},\psi(a_{1},b_{2})=e_{2},\psi(a_{2},b_{1})=e_{3},\psi(a_{2},b_{2})=e_{4},$
where $e_{i}\in\mathbb{R}^{4}$ are the unit base vectors. Directly we have the reward parameters $\theta$ as the rewards.
The behavior policy is $\pi^{b}(a_{1},b_{1})=\pi^{b}(a_{2},b_{2})=1/2$ and dataset is $\mathcal{D}=\{(\tau_{i},\tau^{\prime}_{i},y_{i})\}_{i=1}^{n}$ with
$\tau_{i},\tau^{\prime}_{i}\in\{(a_{1},b_{1}),(a_{2},b_{2}),y_{i}\sim\mathrm{Ber}(\exp(r_{1}(\tau_{i})-r_{1}(\tau^{\prime}_{i})))\}.$
As the dataset covers the Nash equilibrium for both games, with enough samples, we have $U_{\mathcal{D}}(\pi^{*})\leq C$ for any constant $C$. Suppose the output policy of the algorithm is $\pi=(\mu,\nu)$, then $\pi$ is at most 0.5-Nash equilibrium in one of these two games  [^1] See proof for Theorem 3.1 in [Cui and Du, 2022a]. ∎
A.2 Unilateral Policy Coverage
Algorithm 2 Value Estimation
1:Input: Offline dataset $\mathcal{D}$, player index $i$, policy $\pi$.
2:Initialization: $\underline{V}_{H+1,i}^{\pi}(s)=0$.
3:for $h=H,H-1,\cdots,1$ do
4:     $w_{h,i}=[\Sigma_{\mathcal{D},h}^{\mathbb{P}}]^{-1}\sum_{n=1}^{N}\psi(s_{h}^{n},\mathbf{a}_{h}^{n})[\underline{r}_{h,i}(s_{h}^{n},\mathbf{a}_{h}^{n})+\underline{V}^{\pi}_{h+1,i}(s_{h+1}^{n})]$.
5:     $\underline{Q}^{\pi}_{h,i}(\cdot,\cdot)=\max\{\left\langle\psi(\cdot,\cdot),w_{h,i}\right\rangle-C_{\mathbb{P}}[\psi(\cdot,\cdot)^{\top}[\Sigma_{\mathcal{D},h}^{\mathbb{P}}]^{-1}\psi(\cdot,\cdot)]^{1/2},0\}$
6:     $\underline{V}^{\pi}_{h,i}(\cdot)=\mathbb{E}_{a\sim\pi_{h}(\cdot)}\underline{Q}^{\pi}_{h,i}(\cdot,\mathbf{a}).$
7:end for
Algorithm 3 Best Response Estimation
1:Input: Offline dataset $\mathcal{D}$, player index $i$, policy $\pi_{-i}$.
2:Initialization: $\overline{V}_{H+1,i}^{\dagger,\pi_{-i}}(s)=0$.
3:for $h=H,H-1,\cdots,1$ do
4:     $w_{h,i}=[\Sigma_{h,\mathcal{D}}^{\mathbb{P}}]^{-1}\sum_{n=1}^{N}\psi(s_{h}^{n},\mathbf{a}_{h}^{n})[\overline{r}_{h}(s_{h}^{n},\mathbf{a}_{h}^{n})+\overline{V}^{\dagger,\pi_{-i}}_{h+1,i}(s_{h+1}^{n})]$.
5:     $\overline{Q}^{\dagger,\pi_{-i}}_{h,i}(\cdot,\cdot)=\min\{\left\langle\psi(\cdot,\cdot),w_{h,i}\right\rangle+\beta[\psi(\cdot,\cdot)^{\top}[\Sigma_{h,\mathcal{D}}^{\mathbb{P}}]^{-1}\psi(\cdot,\cdot)]^{1/2},H\}$
6:     $\overline{V}^{\dagger,\pi_{-i}}_{h,i}(\cdot)=\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{h,-i}(\cdot)}\overline{Q}^{\dagger,\pi_{-i}}_{h,i}(\cdot,\mathbf{a}).$
7:end for
Algorithm 4 Surrogate Minimization
1:Input: Offline dataset $\mathcal{D}$.
2:Initialization: Algorithm 2 for computing $\underline{V}_{1,i}^{\pi}$ and Algorithm 3 for computing $\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})$.
3:Output: $\pi^{\mathrm{output}}=\operatorname*{argmin}_{\pi}\sum_{i\in[m]}\left[\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-\underline{V}_{1,i}^{\pi}(s)\right]$
We use MLE for reward estimation for each player. For simplicity, we omit the subscript $i$ for player $i$. Note that the reward function can be expressed as $r_{\theta}(\tau)=\sum_{h=1}^{H}r_{h}(s_{h},\mathbf{a}_{h})=\left\langle\psi(\tau),\theta\right\rangle$, where $\theta=[\theta_{1},\theta_{2},\cdots,\theta_{H}]$. At each step $h$, we minimize the NLL loss:
$\widehat{\theta}=\operatorname*{argmin}_{\theta_{h}\leq\sqrt{d},h\in[H]}-\sum_{n=1}^{N}\log\left(\frac{1(y^{n}=1)\exp(r_{\theta}(\tau))}{\exp(r_{\theta}(\tau))+\exp(r_{\theta}(\tau^{\prime}))}+\frac{1(y^{n}=0)\exp(r_{\theta}(\tau^{\prime}))}{\exp(r_{\theta}(\tau))+\exp(r_{\theta}(\tau^{\prime}))}\right).$
By Lemma 5, the confidence region is
$\Theta=\left\{\theta:\left\|\theta-\widehat{\theta}\right\|_{\Sigma_{\mathcal{D}}^{r}+\lambda I}\leq C_{r}=C\sqrt{\frac{dH+\log(1/\delta)}{\lambda^{2}n}+d}\right\}.$
We define the optimistic reward and the pessimistic reward as
$\overline{r}_{h}(s,\mathbf{a}):=\max_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle,\underline{r}_{h}(s,\mathbf{a}):=\min_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle.$
We define the Bellman operator:
$[\mathbb{B}_{h,i}V_{h+1,i}](s,\mathbf{a})=r_{h,i}(s,\mathbf{a})+\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}\mid s,\mathbf{a})V_{h+1,i}(s),\forall i\in[m],h\in[H],s\in\mathcal{S},\mathbf{a}\in\mathcal{A}.$
Lemma 1.
With probability at least $1-\delta$, we have
$r_{h,i}(s,\mathbf{a})-2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\leq\underline{r}_{h,i}(s,\mathbf{a})\leq r_{h,i}(s,\mathbf{a})\leq\overline{r}_{h,i}(s,\mathbf{a})\leq r_{h,i}(s,\mathbf{a})+2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
Proof.
By Lemma 5, with probability at least $1-\delta$, we have $\theta\in\Theta$. Thus we have
$\underline{r}_{h,i}(s,\mathbf{a})=\min_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle\leq r_{h,i}(s,\mathbf{a})\leq\overline{r}_{h,i}(s,\mathbf{a})=\max_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle.$
On the other hand, we have
$\overline{r}_{h,i}(s,\mathbf{a})-r_{h,i}(s,\mathbf{a})$
$=$ $\left\langle\psi(s,\mathbf{a}),\overline{\theta}_{h}\right\rangle-\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle$
$=$ $\left\langle\overline{\psi}_{h}(s,\mathbf{a}),\overline{\theta}-\widehat{\theta}\right\rangle+\left\langle\overline{\psi}_{h}(s,\mathbf{a}),\widehat{\theta}-\theta\right\rangle$
$\leq$ $\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\left\|\overline{\theta}-\widehat{\theta}\right\|_{\Sigma_{\mathcal{D}}^{r}+\lambda I}+\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\left\|\widehat{\theta}-\theta\right\|_{\Sigma_{\mathcal{D}}^{r}+\lambda I}$
$\leq$ $2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
Similarly, we have
$r_{h,i}(s,\mathbf{a})-\underline{r}_{h,i}(s,\mathbf{a})\leq 2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
Lemma 2.
With probability at least $1-\delta$, we have
$\underline{V}^{\pi}_{h,i}(s,\mathbf{a})\leq V^{\pi}_{h,i}(s,\mathbf{a}),\overline{V}^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a})\geq V^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a}),\forall h\in[H],i\in[m],s\in\mathcal{S},\mathbf{a}\in\mathcal{A}.$
Proof.
We prove the arguments by mathematical induction. The arguments hold for step $H+1$ as all quantities are 0. Suppose step $h+1$ holds and we consider step $h$. For the first argument, we have
$\underline{V}_{h,i}^{\pi}(s)=$ $\mathbb{E}_{\mathbf{a}\sim\pi_{h}(s)}\underline{Q}^{\pi}_{h,i}(s,\mathbf{a})$
$\leq$ $\mathbb{E}_{\mathbf{a}\sim\pi_{h}(s)}\left[\mathbb{B}_{h,i}\underline{V}^{\pi}_{h+1,i}(s,\mathbf{a})\right]$ (Lemma 6)
$\leq$ $\mathbb{E}_{\mathbf{a}\sim\pi_{h}(s)}\left[\mathbb{B}_{h,i}V^{\pi}_{h+1,i}(s,\mathbf{a})\right]$ (Lemma 2)
$=$ $V_{h,i}^{\pi}(s).$
For the second argument, we have
$\overline{V}^{\dagger,\pi_{-i}}_{h,i}(s,\mathbf{a})=$ $\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{h,-i}(s)}\overline{Q}^{\dagger,\pi_{-i}}_{h,i}(\cdot,\mathbf{a})$
$\geq$ $\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{h,-i}(s)}\left[\mathbb{B}_{h,i}\overline{V}^{\dagger,\pi_{-i}}_{h+1,i}(s,\mathbf{a})\right]$ (Lemma 6)
$\geq$ $\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{h,-i}(s)}\left[\mathbb{B}_{h,i}V^{\dagger,\pi_{-i}}_{h+1,i}(s,\mathbf{a})\right]$ (Lemma 2)
$=$ $V^{\dagger,\pi_{-i}}_{h,i}(s,\mathbf{a}).$
Lemma 3.
With probability at least $1-\delta$, we have
$V_{1,i}^{\pi}(s_{1})-\underline{V}_{1,i}^{\pi}(s_{1})\leq\mathbb{E}_{\pi}\left[2C_{\mathbb{P}}\sum_{h=1}^{H}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right],$
$\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-V_{1,i}^{\dagger,\pi_{-i}}(s_{1})\leq\mathbb{E}_{\pi_{i}^{\dagger},\pi_{-i}}\left[\sum_{h=1}^{H}2C_{\mathbb{P}}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right].$
Proof.
By Lemma 6, we have
$V_{1,i}^{\pi}(s_{1})-\underline{V}_{1,i}^{\pi}(s_{1})$
$=$ $\mathbb{E}_{\mathbf{a}\sim\pi_{1}(s_{1})}Q^{\pi}_{1,i}(s_{1},\mathbf{a})-\mathbb{E}_{\mathbf{a}\sim\pi_{1}(s_{1})}\underline{Q}^{\pi}_{1,i}(s_{1},\mathbf{a})$
$\leq$ $\mathbb{E}_{\mathbf{a}\sim\pi_{1}(s_{1})}\left[\mathbb{B}_{1,i}V^{\pi}_{2}(s_{1},\mathbf{a})-\mathbb{B}_{1,i}\underline{V}_{2}(s_{1},\mathbf{a})+2C_{\mathbb{P}}\left\|\psi(s_{1},\mathbf{a})\right\|_{[\Sigma_{\mathcal{D},1}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\left\|\overline{\psi}_{1}(s_{1},\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right]$
$=$ $\mathbb{E}_{\mathbf{a}\sim\pi_{1}(s)}\left[V_{2,i}^{\pi}(s_{2})-\underline{V}_{2,i}^{\pi}(s_{2})+2C_{\mathbb{P}}\left\|\psi(s_{1},\mathbf{a})\right\|_{[\Sigma_{\mathcal{D},1}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\left\|\overline{\psi}_{1}(s_{1},\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right]$
$\leq$ $\cdots$
$\leq$ $\mathbb{E}_{\pi}\left[2C_{\mathbb{P}}\sum_{h=1}^{H}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right].$
Similarly, we have
$\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-V_{1,i}^{\dagger,\pi_{-i}}(s_{1})$
$=$ $\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{1,-i}(s_{1})}\overline{Q}_{1,i}(s_{1},\mathbf{a})-\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{1,-i}(s_{1})}Q^{\dagger,\pi_{-i}}_{1,i}(s_{1},\mathbf{a})$
$\leq$ $\mathbb{E}_{\mathbf{a}\sim(\pi_{i}^{\dagger},\pi_{-i})}\left[\overline{Q}_{1,i}(s_{1},\mathbf{a})-Q^{\dagger,\pi_{-i}}_{1,i}(s_{1},\mathbf{a})\right]$
$\leq$ $\mathbb{E}_{\mathbf{a}\sim(\pi_{i}^{\dagger},\pi_{-i})}\left[\mathbb{B}_{1}\overline{V}^{\dagger,\pi_{-i}}_{2}(s_{1},\mathbf{a})-\mathbb{B}_{1}V^{\dagger,\pi_{-i}}_{2}(s_{1},\mathbf{a})+2C_{\mathbb{P}}\left\|\psi(s_{1},\mathbf{a}_{1})\right\|_{[\Sigma_{\mathcal{D}}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\left\|\overline{\psi}_{1}(s_{1},\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right]$
$\leq$ $\cdots$
$\leq$ $\mathbb{E}_{\pi_{i}^{\dagger},\pi_{-i}}\left[\sum_{h=1}^{H}2C_{\mathbb{P}}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right].$
Lemma 4.
Under the event in Lemma 2,
[…]
Proof.
The proof is similar to the proof for Lemma 21 in [Cui and Du, 2022b]. ∎
Theorem 4.
Set $\lambda=1$ for Algorithm 4. With probability $1-\delta$, we have
[…]
where $C_{r}=\widetilde{O}(\sqrt{dH})$ and $C_{\mathbb{P}}=\widetilde{O}(dH)$.
Proof.
By Lemma 3 and Lemma 4, we have
[…]
$\leq$ $\sum_{i\in[m]}\left[\overline{V}_{1,i}^{\dagger,\pi^{*}_{-i}}(s_{1})-\underline{V}_{1,i}^{\pi^{*}}(s)\right]$
$\leq$ $\sum_{i\in[m]}\mathbb{E}_{\pi}\left[2C_{\mathbb{P}}\sum_{h=1}^{H}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right]$
$+\sum_{i\in[m]}\mathbb{E}_{\pi_{i}^{\dagger},\pi_{-i}}\left[\sum_{h=1}^{H}2C_{\mathbb{P}}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right]$
$+\sum_{i\in[m]}\left[V_{1,i}^{\dagger,\pi^{*}_{-i}}(s_{1})-V_{1,i}^{\pi^{*}}(s_{1})\right]$
$\leq$ $\max_{\pi_{i}}4\mathbb{E}_{\pi_{i},\pi^{*}_{-i}}\left[\sum_{h=1}^{H}C_{\mathbb{P}}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}}+C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right],$
where we leverage the fact that $V_{1,i}^{\dagger,\pi^{*}_{-i}}(s_{1})-V_{1,i}^{\pi^{*}}(s_{1})$ for Nash equilibrium $\pi^{*}$. ∎
Intuitively, the proof consists of two main phases: 1) we first reduce the MARL problem to the MARLHF problem, as preference signals can be sampled given the real rewards; 2) we then observe that in MARL problems, a Nash equilibrium is only identifiable when all adjacent actions are represented in the dataset. This observation establishes the necessity of unilateral coverage. The sufficiency of unilateral coverage in MARLHF (Theorems 2, 4) is then derived from its sufficiency in MARL and the reduction from MARL to MARLHF.
A.3 Auxiliary Lemmas
Lemma 5.
(Lemma 3.1 in [Zhu et al., 2023]) With probability at least $1-\delta$, we have
$\left\|\widehat{\theta}-\theta\right\|_{\Sigma_{\mathcal{D}}^{r}+\lambda I}\leq C\sqrt{\frac{d+\log(1/\delta)}{\lambda^{2}}+\lambda B^{2}}.$
Lemma 6.
(Lemma A.1 in [Zhong et al., 2022]) With probability at least $1-\delta$, we have
$0\leq\mathbb{B}_{h}\underline{V}_{h+1,i}(\cdot,\cdot)-\underline{Q}_{h,i}(\cdot,\cdot)\leq 2C_{\mathbb{P}}\left\|\psi(\cdot,\cdot)\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}},$
$0\geq\mathbb{B}_{h}\overline{V}_{h+1,i}(\cdot,\cdot)-\overline{Q}_{h,i}(\cdot,\cdot)\geq-2C_{\mathbb{P}}\left\|\psi(\cdot,\cdot)\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}},$
where $C_{\mathbb{P}}=CdH\sqrt{\log(2dNH/\delta)}$.
Appendix B Experiment Details
B.1 Implementation Details
Model Configurations
The models used in our experiments are designed to effectively handle the complexities of multi-agent environments. Our reward model $r_{\phi}$ is a fully connected neural network, featuring action and observation embedding layers followed by hidden layers. The MADPO agent network uses RNN to output its Q-values, enabling the agent to make informed decisions based on its observations and learned policies. Table 5 lists the main hyperparameters used in our experiments, while other details can be checked in our codebase: https://github.com/NataliaZhang/marlhf.
Hyperparameter Default Value
MSE Loss Coefficient 1
Reference Log Coefficient 1
Prediction Steepness 5
Number of Environment Timesteps 130
Number of Environments 128
Reward Model Type MLP (Spread, Reference), RNN (Tag)
RNN Hidden Size 64 (Tag)
MLP Layer Dimension 128 (Spread), 64 (Reference)
Reward Model Layers 4
Reward Model Epochs 100
Reward Model Learning Rate 1e-3
Reward Model Batch Size 256
IL Learning Rate 1e-3
IL Epochs 100
IL Batch Size 128
VDN Learning Rate 1e-3
VDN Epochs 1e4 (Spread, Reference), 1e6 (Tag)
VDN Batch Size 64
Table 5: Main hyperparameters in experiments
Dataset Configurations
As mentioned in Section 6.2, we collected trajectories in different environments using various policies to ensure a diverse dataset. Our dataset size contains 38400 trajectories in each environment. The number of trajectory pairs is chosen as 10 times the number of trajectories. Preference tags were then generated for these trajectory pairs in the mixed datasets. To adjust the randomness of the preferences, a steepness parameter was introduced as a scalar of the standardized reward. This configuration ensures a comprehensive dataset that can effectively support the evaluation of our methods.
B.2 Tasks Descriptions
MPE is chosen for our experiments due to its versatility and well-established use as a benchmark for MARL algorithms. Among its variety of scenarios, the following three methods are chosen for our experiments:
  • Simple Spread
    • Objective: Group of agents spread out. Each agent aims to occupy a unique landmark while avoiding collisions with other agents.
    • Challenge: There is a potential conflict between the collision penalty and the spreading goal. Any biased policy would pushes the agents away from their targets, leading to suboptimal performance. Successfully balancing these objectives is critical to avoid negative learning outcomes.
  • Simple Tag
    • Objective: Adversaries aim to catch the good agents cooperatively, while good agents aim to avoid being caught.
    • Challenge: The adversaries only gets reward at the timestep of catching a good agent, so recovering the reward distribution across time becomes a challenging work. Note that the original environment is a 1v3 adversary game, and we convert it into a 3-agent cooperative game for better evaluation by fixing the good agent with a MAPPO pretrained policy.
  • Simple Reference
    • Objective: Agents aim to reach target landmarks that are known only to others by communication.
    • Challenge: The requirement for communication increases the complexity of the action space and the dependency among cooperating agents. The performance of agents is affected particularly under unilateral policy conditions, where misaligned communication signals can significantly impact performance.
These tasks provide a robust framework for evaluating the effectiveness and adaptability of our offline MARLHF algorithm in various multi-agent settings. Additionally, they represents the common environments that are sensitive to dataset coverage, where dataset with unilateral policy can easily disrupt cooperation. Therefore, robust approaches are essential to ensure stable performance across different scenarios.
B.3 Scalability Analysis
To evaluate the scalability of our approach, we tested the performance of different methods as the number of agents increased. The experiments were conducted in the Spread-v3 environment, and the test returns per agent were recorded in Table 6.
In our experiments, we observed that as the number of agents increases, convergence times lengthen and the complexity of the problem grows, mirroring the challenges typically encountered in traditional MARL settings. While our current approach manages this scaling without introducing new problems, it does not specifically address the inherent issues of instability and complexity that are well-documented in traditional MARL.
Further work may involve optimizing the algorithms to better handle larger-scale multi-agent environments or exploring alternative methods that maintain high performance even as the agent count increases.
4 agents 5 agents 6 agents 7 agents
Mix-Unilateral -31.13 $\pm$ 0.33 -28.26 $\pm$ 0.43 -26.92 $\pm$ 0.33 -25.48 $\pm$ 0.13
Pure-Expert -31.71 $\pm$ 0.17 -28.80 $\pm$ 0.10 -27.16 $\pm$ 0.39 -26.29 $\pm$ 0.32
Trivial -50.83 -36.92 -28.56 -23.62
Table 6: Test returns per agent in spread-v3 when agent scales. We ran 5 seeds for each dataset and kept all parameters at their default values ($\alpha=1,\beta=1$). Trivial represents test returns where all agents take a random policy, serving as a comparison. As the number of agents scales, the performance of the method generally decrease, and is eventually outperformed by the trivial policy when it reaches 7 agents.
B.4 Ablation Study Details
Figure 3: Reward model training curves on Spread-v3 Diversified dataset. Extra positive MSE regularization results in lower final training loss.
(a) Pure-Expert Spread
(b) Diversified $\alpha=0$ Spread
(c) Diversified Spread
(d) Pure-Expert Tag
(e) Diversified $\alpha=0$ Tag
(f) Diversified Tag
(g) Pure-Expert Reference
(h) Diversified $\alpha=0$ Reference
(i) Diversified Reference
Figure 4: Predicted rewards and ground truth (both standardized) in all environments. Our method with diversified dataset and reward regularization gives predictions that approximate the ground truth the best.
$\alpha$ 0 0.001 0.01 0.1 1 10 100 1000
Spread-v3 0.350 0.345 0.347 0.351 0.361 0.389 0.460 0.603
Tag-v3 0.465 0.431 0.440 0.455 0.484 0.531 0.603 0.676
Reference-v3 0.358 0.356 0.362 0.374 0.393 0.434 0.508 0.623
Table 7: NLL loss over diversified dataset. Appropriate regularization can assist the reward model in learning more effectively, leading to a reduction in NLL loss. Strengthening regularization (larger $\alpha$) sometimes leads to lower NLL loss, indicating better capability of capturing differences in reward signal.
In this section, we explore the effects of specific components of our method, focusing on the influence of MSE regularization and the use of diversified datasets. Our ablation studies collectively underscore the importance of MSE regularization and diversified datasets in enhancing the robustness and accuracy of the reward models within our framework.
The incorporation of MSE regularization plays an important role in improving model stability and convergence. As seen in Figure 3, appropriate regularization leads to lower final training loss, suggesting a more stable learning process. Furthermore, when analyzing reward prediction accuracy in Figure 4, it becomes evident that the combination of diversified datasets and MSE regularization allows the model to better approximate ground truth rewards, indicating enhanced learning capacity.
Additionally, our evaluation of the Negative Log-Likelihood (NLL) loss over diversified datasets reveals that stronger regularization can sometimes lead to a reduction in NLL loss, as listed in Table 7. This implies that the model becomes more capable of capturing nuances in the reward signals as the regularization parameter, $\alpha$, is increased. However, it is also important to balance the strength of regularization, as overly strong regularization could potentially hinder the model’s flexibility in capturing complex reward structures.
The interplay between regularization strength and dataset diversity is critical for achieving optimal model performance in complex multi-agent settings.

### v3

@@
-Multi-Agent Reinforcement Learning from Human Feedback: Data Coverage and Algorithmic Techniques
+Preference-Based Multi-Agent Reinforcement Learning: Data Coverage and Algorithmic Techniques
 Abstract
-We initiate the study of Multi-Agent Reinforcement Learning from Human Feedback (MARLHF), exploring both theoretical foundations and empirical validations.
-We define the task as identifying Nash equilibrium from a preference-only offline dataset in general-sum games, a problem marked by the challenge of sparse feedback signals.
-Our theory establishes the upper complexity bounds for Nash Equilibrium in effective MARLHF, demonstrating that single-policy coverage is inadequate and highlighting the importance of unilateral dataset coverage.
+We initiate the study of Preference-Based Multi-Agent Reinforcement Learning (PbMARL), exploring both theoretical foundations and empirical validations.
+We define the task as identifying the Nash equilibrium from a preference-only offline dataset in general-sum games, a problem marked by the challenge of sparse feedback signals.
+Our theory establishes the upper complexity bounds for Nash Equilibrium in effective PbMARL, demonstrating that single-policy coverage is inadequate and highlighting the importance of unilateral dataset coverage.
 These theoretical insights are verified through comprehensive experiments.
@@
 (1) We propose a Mean Squared Error (MSE) regularization along the time axis to achieve a more uniform reward distribution and improve reward learning outcomes.
-(2) We utilize imitation learning to approximate the reference policy, ensuring stability and effectiveness in training.
-Our findings underscore the multifaceted approach required for MARLHF, paving the way for effective preference-based multi-agent systems.
+(2) We propose an additional penalty based on the distribution of the dataset to incorporate pessimism, improving stability and effectiveness during training.
+Our findings underscore the multifaceted approach required for PbMARL, paving the way for effective preference-based multi-agent systems.
 Keywords multi-agent reinforcement learning  $\cdot$ reinforcement learning from human feedback  $\cdot$ dataset coverage
@@
 In addition, reward design has been a hard problem in multi-agent reinforcement learning (Devlin et al., 2011).
-Thus, it is crucial to further align the multi-agent system with human feedback.
+Thus, it is crucial to further align the multi-agent system from preference feedback.
 We address this problem through both theoretical analysis and empirical experiments.
-Theoretically, we characterize the dataset coverage condition for MARLHF that enables learning the Nash equilibrium, which serves as a favorable policy for each player.
+Theoretically, we characterize the dataset coverage condition for PbMARL that enables learning the Nash equilibrium, which serves as a favorable policy for each player.
 Empirically, we validate our theoretical insights through comprehensive experiments utilizing the proposed algorithmic techniques.
 1.1 Contributions and Technical Novelties
-1. Necessary and Sufficient Dataset Coverage Condition for MARLHF.
+1. Necessary and Sufficient Dataset Coverage Condition for PbMARL.
 In single-agent RLHF, (Zhu et al., 2023) demonstrated that single policy coverage is sufficient for learning the optimal policy.
-However, we prove that this condition no longer holds for MARLHF by providing a counterexample.
+However, we prove that this condition no longer holds for PbMARL by providing a counterexample.
 Instead, we introduce an algorithm that operates under unilateral coverage, a condition derived from offline MARL (Cui and Du, 2022a; Zhong et al., 2022).
@@
 For further details, see Section 4.
-Figure 1: The overall pipeline of offline MARLHF.
-$\mathcal{D}$ is the preference dataset where $\tau$ is a one trajectory and $\mathbf{y}_{i}\in\{1,-1\}^{m}$ indicates which trajectory is preferred by each agent.
+Figure 1: The overall pipeline of offline PbMARL.
+$\mathcal{D}$ is the preference dataset where $\tau_{i},\tau_{i}^{\prime}$ are trajectories and $\mathbf{y}_{i}\in\{1,-1\}^{m}$ indicates which trajectory is preferred by each agent.
 $r_{\phi}$ is the learned reward.
-$\pi_{ref}$ is the learned reference policy using imitation learning.
+$\pi_{b}$ is the learned reference policy using imitation learning.
 2. Algorithmic Techniques for Practical Performance.
-As a foundational exploration into MARLHF research, we focus on employing the simplest learning framework, incorporating only the essential techniques necessary to ensure the approach’s feasibility.
-The framework consists of three key components: 1) leveraging the preference dataset to learn a reward function, 2) mitigating extrapolation errors using a reference policy, and 3) determining the final policy.
+As a foundational exploration into PbMARL research, we focus on employing the simplest learning framework, incorporating only the essential techniques necessary to ensure the approach’s feasibility.
+The framework consists of three key components: 1) leveraging the preference dataset to learn a reward function, 2) mitigating extrapolation errors with pessimism, and 3) determining the final policy.
 Figure 1 provides an overview of the process.
-However, additional algorithmic techniques are needed in order to find a strong policy even if we have a dataset with good coverage according to the theory.
+However, additional algorithmic techniques are required to identify a robust policy, even when the dataset demonstrates good coverage according to our theoretical insights.
   • Reward regularization.
@@
 This regularization helps to prevent the model from accumulating reward signals solely at the final time step or relying on reward-irrelevant observation patterns, which could otherwise result in the complete failure in producing meaningful predictions.
-  • Imitation learning policy as a reference.
-We choose to follow the standard RLHF work (Wang et al., 2023b), where the extrapolation problem was alleviated by adding an extra KL reward to the final optimization objective.
-Since we do not have access to the actual reference model, we implement imitation learning across the entire dataset to approximate it.
-The final policy is trained using a DQN-based Value Decomposition Network (VDN) (Mnih et al., 2013; Sunehag et al., 2017).
-Our ablation study demonstrates the effectiveness of selecting an appropriate KL reward coefficient (see Table 3).
+  • Dataset Distribution-Based Pessimism.
+To mitigate the extrapolation error in offline RL, we add an extra reward term based on the density of a certain state-action pair in the dataset to implement pessimism.
+In our approach, an imitation learning agent is trained to model the density function.
+The final policy is then trained using a DQN-based Value Decomposition Network (VDN) (Mnih et al., 2013; Sunehag et al., 2017).
+Our ablation study demonstrates the critical role of appropriately tuning the reward coefficient to ensure training stability and performance (see Table 4).
 3. Experiment Results.
 Our experiments, following the pipeline described above, confirm the theoretical necessity of unilateral coverage.
-We conducted comprehensive ablation studies on three cooperative Multi-Agent Particle Environment (MPE) scenarios (Mordatch and Abbeel, 2017): Spread-v3, Tag-v3, and Reference-v3.
-These studies focused on the hyperparameter selection for the reward regularization coefficient, KL reward coefficient, and dataset diversity.
-The empirical results (Table 2) demonstrate that: 1) simply adding trivial trajectories to expert demonstrations can enhance performance, 2) unilateral datasets are advantageous, and 3) dataset diversity contributes to lower variance.
+We performed extensive ablation studies across three Multi-Agent Particle Environment (MPE) scenarios—Spread-v3, Tag-v3, and Reference-v3 (Mordatch and Abbeel, 2017)—as well as the popular Overcooked environment (Carroll et al., 2020).
+These studies focused on the hyperparameter selection for the reward regularization coefficient $\alpha$, pessimism coefficient $\beta$, and dataset diversity.
+The empirical results (Table 2) demonstrate that: 1) augmenting expert demonstrations with trivial trajectories significantly improves performance, 2) unilateral datasets are advantageous, and 3) dataset diversity contributes to lower variance.
 Our ablation experiments underscore the effectiveness of the proposed algorithmic techniques.
-Additionally, we tested the sensitivity of the hyperparameters used in MSE and KL regularization, and we introduce a principled standardization technique that can efficiently tune hyperparameters across all environments and datasets.
+Additionally, we introduced a principled standardization technique that efficiently tunes hyperparameters across all environments and datasets.
 2 Related Works
 Reinforcement Learning from Human Feedback (RLHF).
-RLHF (or preference-based RL) plays a pivotal role in alignment with various tasks such as video games (Warnell et al., 2018; Brown et al., 2019), robotics (Jain et al., 2013; Kupcsik et al., 2016; Christiano et al., 2023; Shin et al., 2023), image augmentation (Metcalf et al., 2024), and large language models (Ziegler et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022; Stiennon et al., 2022; Bai et al., 2022; Glaese et al., 2022; Ganguli et al., 2022; Ouyang et al., 2022).
+RLHF, or preference-based RL (PbRL), plays a pivotal role in alignment with various tasks such as video games (Warnell et al., 2018; Brown et al., 2019), robotics (Jain et al., 2013; Kupcsik et al., 2016; Christiano et al., 2023; Shin et al., 2023), image augmentation (Metcalf et al., 2024), and large language models (Ziegler et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022; Stiennon et al., 2022; Bai et al., 2022; Glaese et al., 2022; Ganguli et al., 2022; Ouyang et al., 2022).
 Additionally, a body of work focuses on the reward models behind preference data (Sadigh et al., 2017; Bıyık and Sadigh, 2018; Gao et al., 2022; Hejna and Sadigh, 2023).
-Recently, direct preference optimization (DPO, Rafailov et al.
+Recent works like VIPO (Cen et al., 2024) incorporates uncertainty-aware regularization into the reward model, while (Liu et al., 2024) address over-optimization using adversarial regularization.
+Direct preference optimization (DPO, Rafailov et al.
 (2023)) and its variants (Azar et al., 2023; Rafailov et al., 2024) approach RLHF without directly handling the reward model.
-Theoretical studies have also explored guarantees, such as sample complexity and regret, and the limitations of certain RLHF algorithms (Novoseller et al., 2020; Xu et al., 2020; Pacchiano et al., 2023; Chen et al., 2022; Razin et al., 2023; Zhu et al., 2024a; Wang et al., 2023b; Xiong et al., 2024; Zhu et al., 2024b).
+Theoretical studies have also explored guarantees, such as sample complexity and regret, and the limitations of certain RLHF algorithms (Novoseller et al., 2020; Xu et al., 2020; Pacchiano et al., 2023; Chen et al., 2022; Razin et al., 2023; Zhu et al., 2024a; Wang et al., 2023c; Xiong et al., 2024; Zhu et al., 2024b).
 Offline Reinforcement Learning.
@@
 (2022); Xiong et al.
-(2023); Li et al.
+(2023b); Li et al.
+(2024); Xie et al.
+(2023); Mete et al.
 (2021).
@@
 Prominent algorithms in MARL include IQL (Tan, 2003), MADDPG (Lowe et al., 2020), COMA (Foerster et al., 2017), MAPPO (Yu et al., 2022), VDN (Sunehag et al., 2017), and QMIX (Rashid et al., 2018).
+Theoretical research has made great process in reducing the sample complexity (Wang et al., 2023b; Xiong et al., 2023a).
 Offline MARL.
@@
 (2023); Shi et al.
+(2023); Zhong et al.
 (2022).
@@
 In this section, we study the dataset coverage assumptions for offline MARLHF.
-For offline single-agent RLHF, (Zhu et al., 2023; Zhan et al., 2023) shows that single policy coverage is sufficient for learning the optimal policy.
+For offline single-agent RLHF, Zhu et al.
+(2023); Zhan et al.
+(2023) show that single policy coverage is sufficient for learning the optimal policy.
 However, we prove that this assumption is insufficient in the multi-agent setting by constructing an counterexample.
@@
 5 Algorithmic Techniques for Practical Performance
-In Section 4, we provided a theoretical characterization of the dataset requirements for MARLHF.
+In Section 4, we provided a theoretical characterization of the dataset requirements for PbMARL.
 However, the algorithm used in Theorem 2 is not computationally efficient.
-In this section, we propose a practical algorithm for MARLHF and validate our theoretical findings through experiments.
+In this section, we propose a practical algorithm for PbMARL and validate our theoretical findings through experiments.
 5.1 High-level Methodology
-Our MARLHF pipeline consists of two phases: In the first step, we train a reward prediction model $\phi$ and approximate the behavior policy $\pi^{b}$ using imitation learning; in the second step, we then apply an MARL algorithm to maximize a combination of the KL-divergence-based reward and standardized predicted reward $r_{\phi}$, ultimately deriving the final policy $\pi_{\mathbf{w}}$.
-Step 1: Reward Training and Reference Approximation.
+Our PbMARL pipeline consists of two phases: In the first step, we train a reward prediction model $\phi$ and approximate the behavior policy $\pi_{b}$ using imitation learning; in the second step, we then apply an MARL algorithm to maximize a combination of the KL-divergence-based reward and standardized predicted reward $r_{\phi}$, ultimately deriving the final policy $\pi_{\mathbf{w}}$.
+Step 1: Reward Training and Dataset Modeling.
 Given the preference signals of trajectories, we use neural networks to predict step-wise rewards $r_{\phi}(s_{h},a_{h})$ for each agent, minimizing the loss defined in (1).
 The objective is to map $(s,a_{i})$-pairs to reward values such that the team returns align with the preference signals.
-At the same time, in order to utilize KL reward $-\log\frac{\pi_{\mathbf{w}}(s,a)}{\pi^{b}(s,a)}$ to cope with the extrapolation error in offline learning, an imitation learner $\pi_{\text{ref}}$ is trained over the entire dataset to model the behavior policy $\pi^{b}$.
+At the same time, in order to utilize distribution-based penalty term $\log{\pi_{b}(s,a)}$ to cope with the extrapolation error in offline learning, an imitation learner is trained over the entire dataset to model the behavior policy $\pi_{b}$.
 Step 2: Offline MARL.
 Although in this work, VDN is chosen as the MARL oracle, it should be noted that other MARL architectures are also applicable.
-With the reward model $r_{\phi}$ and the approximated reference policy $\pi_{\text{ref}}$ learned in the first step, we are now able to construct a virtual step-wise reward for each agent.
+With the reward model $r_{\phi}$ and the approximated dataset distribution learned in Step 1, we are now able to construct a virtual step-wise reward for each agent.
 The agents are then trained to maximize the target defined in (3).
@@
 Section 6.3).
-5.3 Imitation Learning Policy as Reference
+5.3 Dataset Distribution-Based Pessimism
 There are various methods to mitigate the over-extrapolation errors in offline RL (Peng et al., 2019; Nair et al., 2021), including conservative loss over the Q-function (Kumar et al., 2020) and directly restricting the learned policy actions to those within within the dataset (Fujimoto et al., 2019).
-For simplicity and consistency with the former RLHF framework (Ouyang et al., 2022), we adopt the same per-step KL reward for enforcing restrictions between $\pi^{b}$ and $\pi_{\mathbf{w}}$.
-In many scenarios, a direct reference to the behavior policy is not accessible.
-As an alternative, imitation learning is used to estimate the reference policy, $\pi_{\text{ref}}\approx\pi^{b}$.
-To stabilize training, we standardize $r_{\phi}$ over $\mathcal{D}$ before combining it with the KL reward to make it comparable:
-$\text{objective}(\mathbf{w})=\mathbb{E}_{\tau\sim\pi_{\mathbf{w}}}\left[\sum_{h=1}^{H}r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)-\beta\log\frac{\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})}{\pi_{\text{ref}}(s_{h},\mathbf{a}_{h})}\right],$ (3)
-where $\beta$ is the KL reward coefficient, set to be (1, 1, 3) in Spread-v3, Reference-3 and Tag-v3 respectively, and the standardized reward $r_{\text{std}}$ is defined as:
+We add a per-step dataset-based penalty term, $\log{\pi_{b}(s,\mathbf{a})}$, as pessimism towards less explored states.
+Imitation learning is utilized to estimate the behavior policy $\pi_{b}$ from the dataset distribution.
+To stabilize training, we standardize predicted reward $r_{\phi}$ over $\mathcal{D}$ before combining it with the penalty term to make them comparable:
+$\text{objective}(\mathbf{w})=\mathbb{E}_{\tau\sim\pi_{\mathbf{w}}}\left[\sum_{h=1}^{H}r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)+\beta\log{\pi_{b}(s_{h},\mathbf{a}_{h})}\right],$ (3)
+[…]
 $r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)=\sum_{i=1}^{m}\frac{r_{\phi}(s_{h},a_{h,i})-\mathbb{E}_{\mathcal{D}}({r}_{\phi})}{\sqrt{\text{Var}_{\mathcal{D}}(r_{\phi})}}.$ (4)
-Notice that the explicit expression of $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ only exists in methods with an actor model, which leads to certain planning-based algorithms, such as Q-learning methods, being unable to directly use Equation 3.
-To enhance the compatibility of our pipeline, we substitute $1/|A|$ for $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ in such cases.
-Although this substitution formally removes the KL reward, in algorithms like Q-learning that output deterministic policies, the replaced reward is equivalent to encouraging the agent to match its deterministic choices with the actions most preferred by the reference policy.
-In other words, it seeks the closet element in the deterministic policy family to the reference policy, which is a degradation of the KL reward restricted in the domain of deterministic policy.
+Intuitively, the penalty term $\log{\pi_{b}(s_{h},\mathbf{a}_{h})}$ discourages the agents from deviating from the most preferred actions in the dataset.
 The effectiveness of this method is validated in the ablation study (cf.
@@
 We design a series of experiments to validate our theories and methods in common general-sum games.
-Specifically, we first use VDN Sunehag et al.
-(2017) to train expert agents, and take intermediate checkpoints as rookie agents.
-Then, we use these agents to collect datasets and use the Broadley-Terry model over standardized returns to simulate human preference.
+Specifically, we first use online RL algorithms to train expert agents, and take intermediate checkpoints as rookie agents.
+Then, we use these agents to collect datasets and use the Bradley-Terry model over standardized returns to simulate human preference.
 Experiments are carried out to verify the efficiency of our approach with unilateral policy dataset coverage (in Theorem 2) while single policy coverage is insufficient (stated in Theorem 1).
-We also design ablation studies to showcase the importance of our methods, particularly focusing on reward structure regularization and imitation poclies as reference models.
+We also design ablation studies to showcase the importance of our methods, particularly focusing on reward regularization and dataset distribution-based pessimism.
 6.1 Environments
-Our experiments involved 3 Multi-Agent Particle Environments (MPE) implemented with JaxMARL codebase (Rutherford et al., 2023).
-Specifically, we utilized the Spread, Tag, and Reference scenarios, which covers three different levels of communication requirements.
+Our experiments involved three Multi-Agent Particle Environments (MPE), including Spread-v3, Tag-v3 and Reference-v3, and Overcooked environment implemented with JaxMARL codebase (Rutherford et al., 2023).
 Spread-v3 contains a group of agents and target landmarks, where the objective is to cover as many landmarks as possible while avoiding collisions.
@@
 To ensure a fair comparison of different predator cooperation policies, we fixed a pretrained prey agent.
-Reference-v3 involves two agents and potential landmarks, where the agents need to find each one’s target landmark to receive a high reward.
-The target landmark of each agent is only known by the other agent at first, necessitating communication before they can move toward the correct landmark.
+Reference-v3 involves two agents and three potential landmarks, where the agents need to find each one’s target landmark to receive a high reward.
+The target landmark of each agent is only known by the other agent at first.
+Overcooked involves two agents moving and operating objects in a gridworld.
 A more detailed description of the tasks and their associated challenges is provided in Appendix B.2.
@@
 To study the influence of diversity of dataset, we manually designed 4 kinds of mixed joint behavior polices, and change their ratios to form different datasets.
-  • Expert policy: $n$ expert agents. Trained with VDN till convergence.
-  • Rookie policy: $n$ rookie agents. Trained with VDN with early stop.
+  • Expert policy: $n$ expert agents. Trained with online RL algorithms till convergence.
+  • Rookie policy: $n$ rookie agents. Trained with online RL algorithms with early stop.
   • Trivial policy: $n$ random agents. All actions are uniformly sampled from the action space.
@@
 […]
-The experimental results are presented in Table 2.
-We can clearly observe that trivial comparisons (Mix-Expert vs.
-Pure-Expert) play a crucial role in learning an appropriate reward.
-A dataset containing only expert policies proves hard for the reward model to learn from the very beginning.
-In more challenging tasks, as reflected by higher MSE, the importance of unilateral coverage and diversity becomes more pronounced.
-Suboptimal cooperative agent actions assist the reward model in identifying the Nash equilibrium, thus highlighting key reward signals more effectively, as suggested by Theorem 2.
-Additionally, we found that in a diversified dataset, the variance across all training runs was lower compared to other datasets.
-Beyond test returns, we also visualized the reward predictions and utilized the MSE between standardized reward predictions and ground truth as a measure of prediction quality.
-Although this MSE is not necessarily directly correlated with final performance—since accurate predictions do not need to match the shape of the ground truth—it can serve as a useful empirical metric.
+Due to the inherent limitations of offline reinforcement learning (RL) in action selection dictated by the dataset, the effectiveness of learning is often strongly correlated with dataset quality, i.e.
+the level of expertise demonstrated in the dataset.
+However, the results in preference-based MARL experiments partially diverge from this conventional conclusion.
+While the quality of the dataset remains critical, experiments on Reference-v3 and Overcooked (Table 2) indicate that diversity and unilateral data can significantly enhance the performance of the reward model, thereby facilitating learning.
+The main experimental results are presented in Table 2 and Table 3.Among all the experiments, apart from the experiments on Tag-v3, where the high operational precision requirements make data quality more critical than diversity, the other three environments validate our conclusions across all algorithms.
 Expert Unilateral Rookie Trivial
@@
 Pure-Expert 4 0 0 0
-Table 1: Final datasets mixed with various ratios. The overall dataset size is kept to 38400 trajectories. More details list in Appendix B.1.
-Spread-v3 Tag-v3 Reference-v3 Spread-v3 Tag-v3 Reference-v3
-Return Return Return MSE MSE MSE
-Diversified -21.63 $\pm$ 0.51 28.94 $\pm$ 1.04 -19.59 $\pm$ 0.55 0.51 1.04 0.55
-Mix-Unilateral -20.98 $\pm$ 0.56 32.04 $\pm$ 1.12 -20.60 $\pm$ 1.15 0.56 1.12 1.15
-Mix-Expert -21.11 $\pm$ 1.16 33.01 $\pm$ 0.70 -20.38 $\pm$ 1.06 1.16 0.70 1.06
-Pure-Expert -21.45 $\pm$ 1.01 40.14 $\pm$ 0.75 -28.19 $\pm$ 0.92 1.01 0.75 0.92
-Table 2: Test returns and the mean squared error (MSE) between the standardized predicted rewards and the standardized ground truth rewards.
-Detailed descriptions of the datasets are provided in Section 6.2.
-In environments with smoother reward distributions, such as Spread-v3 and Reference-v3, dataset diversity primarily reduces variances.
-However, in more challenging environments, such as Tag-v3, dataset diversity plays a substantially more significant role.
-6.3 Experiments for Reward Regularization
-In Figure 2, we examined the effectiveness of our proposed reward regularization technique.
-Figure 2(a) demonstrates that without regularization, the learned rewards tend to be sparse and spiky compared to the ground truth rewards.
-The predicted rewards in (b1) and (b3) were learned using our reward regularization, while those in (b2) were learned without it ($\alpha=0$).
-Our results indicate that the reward regularization technique produces much smoother reward functions.
-We also quantitatively compare learned rewards with and without using reward regularization in finding the policy (cf.
-the first and the last column of Table 3).
-We also observe that the rewards often exhibit temporal continuity, which can create greater discrepancies with the sparse, pulse-like ground truth.
-Notably, we found that adding stronger regularization does not necessarily lead to underfitting of the reward model; in some cases, it even helps the model converge to a lower training loss.
-Detailed parameters and experimental results are provided in the appendix (cf.
-Table 7).
-We attribute this to the role of regularization in preventing the model from overly relying on shortcuts.
+Table 1: Final datasets mixed with various ratios. The overall dataset size is kept to 38400 trajectories for MPE, and 960 trajectories for Overcooked. (cf. Appendix B.1)
+Algorithm Dataset Spread-v3 Tag-v3 Reference-v3 Overcooked
+VDN with Diversified -21.16 $\pm$ 0.54 29.28 $\pm$ 1.08 -18.89 $\pm$ 0.60 238.89 $\pm$ 3.50
+Pessimism Penalty Mix-Unilateral -21.03 $\pm$ 0.44 36.65 $\pm$ 0.70 -18.80 $\pm$ 0.63 221.80 $\pm$ 26.66
+Mix-Expert -20.98 $\pm$ 0.54 35.96 $\pm$ 0.86 -18.80 $\pm$ 0.44 35.26 $\pm$ 55.19
+Pure-Expert -21.01 $\pm$ 0.57 39.55 $\pm$ 0.77 -28.97 $\pm$ 2.89 3.36 $\pm$ 7.19
+Table 2: In the simplest environment, Spread-v3, different dataset gives similar performance.
+In Tag-v3 environment, where precise actions are required, the quality of the dataset (proportion of expert demonstration) is more important than diversity.
+In contrast, in Overcooked environment, which focuses on strategy learning and demands less on precision, dataset diversity contributes to improved stability, with Unilateral playing a particularly critical role.
+In the Reference-v3 environment, which balances the need for precision and strategic, the importance of both factors is more balanced, but non-expert data is still necessary.
 (a) Mix-Expert $\alpha=0$ (spread-v3)
@@
 Our method with diversified dataset (b3) gives predictions that approximate the ground truth well.
+Algorithm Dataset Spread-v3 Reference-v3 Overcooked
+MAIQL Diversified -25.33 $\pm$ 1.40 -22.15 $\pm$ 0.55 16.59 $\pm$ 11.22
+Mix-Unilateral -23.25 $\pm$ 1.06 -23.22 $\pm$ 1.37 0.00 $\pm$ 0.00
+Mix-Expert -23.26 $\pm$ 0.90 -24.21 $\pm$ 1.60 0.00 $\pm$ 0.00
+Pure-Expert -26.01 $\pm$ 1.53 -29.47 $\pm$ 1.65 0.00 $\pm$ 0.00
+MABCQ Diversified -20.02 $\pm$ 0.64 -17.64 $\pm$ 0.43 239.34 $\pm$ 1.67
+Mix-Unilateral -19.47 $\pm$ 0.33 -17.64 $\pm$ 1.11 215.01 $\pm$ 65.43
+Mix-Expert -19.42 $\pm$ 0.17 -17.88 $\pm$ 0.78 50.32 $\pm$ 82.82
+Pure-Expert -20.56 $\pm$ 0.38 -25.90 $\pm$ 1.11 1.14 $\pm$ 3.46
+Table 3: Test returns of MAIQL and MABCQ.
+In the experimental results, we can observe a clear preference toward more diversified datasets.
+Compared to our method and BCQ, which directly calculate $\max_{a}Q$ for Bellman updates, IQL employs expectile regression to estimate it.
+So MAIQL demands higher accuracy of the reward model.
+Consequently, the performance improvements brought by dataset diversity are also more pronounced in MAIQL experiments.
 $\beta=0$ $\beta=0.1$ $\beta=1$ $\beta=10$ $\beta=100$ $\alpha=0$
@@
 Reference-v3 -19.69 $\pm$ 0.36 -19.37 $\pm$ 0.53 -18.89 $\pm$ 0.78 -18.33 $\pm$ 0.42 -18.54 $\pm$ 0.46 -21.86 $\pm$ 0.73
-Table 3: Comparison of test return per agent with different parameters.
-Standard pipeline take KL reward coefficient $\beta=1$ for Spread-v3, Reference-v3 and $\beta=10$ for Tag-v3, and the MSE reward regularization coefficient $\alpha$ is set to the optimal value for fixed $\beta$.
-All the agents are trained on Diversified Dataset across 5 random seeds.
-This experiment shows the importance using reward regularization and reference policy.
-6.4 Experiments for Imitation Learning
-Intuitively, the choice of the KL coefficient $\beta$ depends on the quality of the reward model $\phi$ and the dataset.
-As shown in Table 3, simpler environments (e.g., Spread and Reference) require less KL restriction, while harder environments (e.g., Tag) need stronger restriction to avoid suffering from the extrapolation problem.
-A large $\beta$ can amplify the approximation error of imitation learning, potentially overshadowing the reward signal, whereas a small $\beta$ may fail to introduce enough pessimism to effectively mitigate extrapolation errors.
-In environments with relatively simple reward structures, such as Spread-v3 and Reference-v3, the standard choice of $\beta=1$ is near optimal.
-However, in more complex scenarios where the reward is harder to derive, increasing $\beta$ to 3 yields better performance.
+Overcooked 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 149.53 $\pm$ 86.74 238.89 $\pm$ 3.50 240 $\pm$ 0.00 240 $\pm$ 0.00
+Table 4: Comparison of test return with different hyperparameters.
+Standard pipeline take pessimism coefficient $\beta=1$ for Spread-v3, Reference-v3 and $\beta=10$ for Tag-v3, Overcooked, and the MSE reward regularization coefficient $\alpha$ is set to the optimal value for fixed $\beta$.
+All the agents are trained on Diversified Dataset across 10 random seeds.
+Results show that larger $\beta$ always gives better performance and a proper positive $\alpha$ can improve performance.
+6.3 Other Ablation Studies
+Reward regularization
+In Figure 2, we examined the effectiveness of our proposed reward regularization technique.
+Figure 2(a) demonstrates that without regularization, the learned rewards tend to be sparse and spiky compared to the ground truth rewards.
+We also observe that the rewards often exhibit temporal continuity, which can create greater discrepancies with the sparse, pulse-like ground truth.
+Notably, we found that adding stronger regularization does not necessarily lead to underfitting of the reward model; in some cases, it even helps the model converge to a lower training loss.
+Detailed parameters and experimental results are provided in the appendix (cf.
+Table 8).
+We attribute this to the role of regularization in preventing the model from overly relying on shortcuts.
+Pessimism coefficient
+Due to the clipping in Equation 3, excessively large $\beta$ values will not dominate the entire reward function.
+As a result, larger $\beta$ values almost never degrade the agent’s performance in our experiments (Table 4).
+This allows us to increase $\beta$ with relative confidence.
+Therefore, we generally recommend setting $\beta$ to a value between 10 and 100 for optimal performances.
+Scalability
+We also tested the scalability on Spread-v3.
+While our current approach manages the scaling of agents without introducing new problems, it does not specifically address the inherent issues of instability and complexity that are well-documented in traditional MARL (cf.
+Appendix B.3).
 7 Discussion
-In this paper, we proposed dedicated algorithmic techniques for offline MARLHF and provided theoretical justification for the unilateral dataset coverage condition.
-We believe our work is a significant step towards systematically studying MARLHF and offers a foundational framework for future research in this area.
+In this paper, we proposed dedicated algorithmic techniques for offline PbMARL and provided theoretical justification for the unilateral dataset coverage condition.
+We believe our work is a significant step towards systematically studying PbMARL and offers a foundational framework for future research in this area.
 The flexibility of our framework allows for application across a wide range of general games, and our empirical results validate the effectiveness of our proposed methods in various scenarios.
 Looking ahead, there is significant potential to extend this work to more complex, real-world scenarios, particularly by integrating Large Language Models (LLMs) into multi-agent systems.
-Future research will focus on fine-tuning and aligning LLMs within MARLHF, addressing challenges such as increased complexity and the design of effective reward structures.
+Future research will focus on fine-tuning and aligning LLMs within PbMARL, addressing challenges such as increased complexity and the design of effective reward structures.
 Appendix A Missing Proofs in Section 4
@@
 3:Output: $\pi^{\mathrm{output}}=\operatorname*{argmin}_{\pi}\sum_{i\in[m]}\left[\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-\underline{V}_{1,i}^{\pi}(s)\right]$
-We use MLE for reward estimation for each player.
+In our framework, we utilize Maximum Likelihood Estimation (MLE) to estimate the reward function for each player.
 For simplicity, we omit the subscript $i$ for player $i$.
-Note that the reward function can be expressed as $r_{\theta}(\tau)=\sum_{h=1}^{H}r_{h}(s_{h},\mathbf{a}_{h})=\left\langle\psi(\tau),\theta\right\rangle$, where $\theta=[\theta_{1},\theta_{2},\cdots,\theta_{H}]$.
+Note that the reward function can be expressed as $r_{\theta}(\tau)=\sum_{h=1}^{H}r_{h}(s_{h},\mathbf{a}_{h})=\left\langle\psi(\tau),\theta\right\rangle$, where $\theta=[\theta_{1},\theta_{2},\cdots,\theta_{H}]$ represents the parameters we aim to optimize.
 At each step $h$, we minimize the NLL loss:
 $\widehat{\theta}=\operatorname*{argmin}_{\theta_{h}\leq\sqrt{d},h\in[H]}-\sum_{n=1}^{N}\log\left(\frac{1(y^{n}=1)\exp(r_{\theta}(\tau))}{\exp(r_{\theta}(\tau))+\exp(r_{\theta}(\tau^{\prime}))}+\frac{1(y^{n}=0)\exp(r_{\theta}(\tau^{\prime}))}{\exp(r_{\theta}(\tau))+\exp(r_{\theta}(\tau^{\prime}))}\right).$
-By Lemma 5, the confidence region is
+This optimization problem helps in learning a reward function that aligns well with the observed data.
+By Lemma 5, we establish a confidence region for the estimated parameters $\theta$:
 $\Theta=\left\{\theta:\left\|\theta-\widehat{\theta}\right\|_{\Sigma_{\mathcal{D}}^{r}+\lambda I}\leq C_{r}=C\sqrt{\frac{dH+\log(1/\delta)}{\lambda^{2}n}+d}\right\}.$
-We define the optimistic reward and the pessimistic reward as
+We define the optimistic reward and the pessimistic reward as follows:
 $\overline{r}_{h}(s,\mathbf{a}):=\max_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle,\underline{r}_{h}(s,\mathbf{a}):=\min_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle.$
@@
 $r_{h,i}(s,\mathbf{a})-2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\leq\underline{r}_{h,i}(s,\mathbf{a})\leq r_{h,i}(s,\mathbf{a})\leq\overline{r}_{h,i}(s,\mathbf{a})\leq r_{h,i}(s,\mathbf{a})+2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
+This lemma establishes bounds on the reward function $r_{h,i}(s,\mathbf{a})$.
+Specifically, the optimistic reward and pessimistic reward are under constrains at a high probability, up to a margin of error determined by $C_{r}$ and the norms of the feature representation $\overline{\psi}_{h}(s,\mathbf{a})$.
 Proof.
-By Lemma 5, with probability at least $1-\delta$, we have $\theta\in\Theta$. Thus we have
+We begin by referencing Lemma 5, which establishes that with probability at least $1-\delta$, the estimated parameters $\theta$ reside within the confidence region $\Theta$.
+This allows us to assert the following relationships:
 $\underline{r}_{h,i}(s,\mathbf{a})=\min_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle\leq r_{h,i}(s,\mathbf{a})\leq\overline{r}_{h,i}(s,\mathbf{a})=\max_{\theta\in\Theta}\left\langle\psi(s,\mathbf{a}),\theta_{h}\right\rangle.$
-On the other hand, we have
+Next, we quantify the deviation between the optimistic reward and the true reward:
 $\overline{r}_{h,i}(s,\mathbf{a})-r_{h,i}(s,\mathbf{a})$
@@
 $\leq$ $2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
-Similarly, we have
+A similar argument can be applied to show that the pessimistic reward is also bounded:
 $r_{h,i}(s,\mathbf{a})-\underline{r}_{h,i}(s,\mathbf{a})\leq 2C_{r}\left\|\overline{\psi}_{h}(s,\mathbf{a})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}.$
 Lemma 2.
-With probability at least $1-\delta$, we have
-$\underline{V}^{\pi}_{h,i}(s,\mathbf{a})\leq V^{\pi}_{h,i}(s,\mathbf{a}),\overline{V}^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a})\geq V^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a}),\forall h\in[H],i\in[m],s\in\mathcal{S},\mathbf{a}\in\mathcal{A}.$
+With probability at least $1-\delta$, the following bounds hold for all time steps $h\in[H]$, for each agent $i\in[m]$, for all states $s\in\mathcal{S}$, and actions $a\in\mathcal{A}$:
+$\underline{V}^{\pi}_{h,i}(s,a)$ $\leq V^{\pi}_{h,i}(s,a),$
+$\overline{V}^{\dagger,\pi_{-i}}_{h}(s,a)$ $\geq V^{\dagger,\pi_{-i}}_{h}(s,a).$
 Proof.
-We prove the arguments by mathematical induction. The arguments hold for step $H+1$ as all quantities are 0. Suppose step $h+1$ holds and we consider step $h$. For the first argument, we have
+We prove the statements by mathematical induction.
+The base case for step $H+1$ holds trivially, as all quantities are zero at this final step.
+Suppose step $h+1$ holds and we consider step $h$.
+We will show they also hold for step $h$.
+For the first argument of pessimistic value, we have
 $\underline{V}_{h,i}^{\pi}(s)=$ $\mathbb{E}_{\mathbf{a}\sim\pi_{h}(s)}\underline{Q}^{\pi}_{h,i}(s,\mathbf{a})$
@@
 $=$ $V_{h,i}^{\pi}(s).$
-For the second argument, we have
+For the second argument of optimistic value, we have
 $\overline{V}^{\dagger,\pi_{-i}}_{h,i}(s,\mathbf{a})=$ $\max_{a_{i}\in\mathcal{A}_{i}}\mathbb{E}_{\mathbf{a}_{-i}\sim\pi_{h,-i}(s)}\overline{Q}^{\dagger,\pi_{-i}}_{h,i}(\cdot,\mathbf{a})$
@@
 $=$ $V^{\dagger,\pi_{-i}}_{h,i}(s,\mathbf{a}).$
+These complete the induction step and shows the lemma holds for all steps. ∎
 Lemma 3.
@@
 $\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-V_{1,i}^{\dagger,\pi_{-i}}(s_{1})\leq\mathbb{E}_{\pi_{i}^{\dagger},\pi_{-i}}\left[\sum_{h=1}^{H}2C_{\mathbb{P}}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right].$
+This lemma establishes bounds on the difference between the expected value of the policy and its lower or upper estimates.
+Using properties of the Bellman operator, we can show that the difference between $V_{1,i}^{\pi}(s_{1})$ and $\underline{V}_{1,i}^{\pi}(s_{1})$ is controlled by the expected norms of the feature representations $\psi$ and $\overline{\psi}$, scaled by constants $C_{\mathcal{P}}$ and $C_{r}$.
 Proof.
-By Lemma 6, we have
+We prove each bound separately, utilizing the results from Lemma 6.
+For the first bound of pessimistic value estimation, we analyze the expected action-value $Q^{\pi}_{1,i}(s_{1},\mathbf{a})$ and its lower estimate, applying the Bellman operator iteratively. We have
 $V_{1,i}^{\pi}(s_{1})-\underline{V}_{1,i}^{\pi}(s_{1})$
@@
 $\leq$ $\mathbb{E}_{\pi}\left[2C_{\mathbb{P}}\sum_{h=1}^{H}\left\|\psi(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D},h}^{\mathbb{P}}+\lambda I]^{-1}}+2C_{r}\sum_{h=1}^{H}\left\|\overline{\psi}_{h}(s_{h},\mathbf{a}_{h})\right\|_{[\Sigma_{\mathcal{D}}^{r}+\lambda I]^{-1}}\right].$
-Similarly, we have
+Similarly, for optimistic value estimation, we have
 $\overline{V}_{1,i}^{\dagger,\pi_{-i}}(s_{1})-V_{1,i}^{\dagger,\pi_{-i}}(s_{1})$
@@
 Proof.
-The proof is similar to the proof for Lemma 21 in [Cui and Du, 2022b]. ∎
+[…]
+$\leq\sum_{i\in[m]}\left[\overline{V}_{1,i}^{\dagger,\pi^{*}_{-i}}(s_{1})-\underline{V}_{1,i}^{\pi}(s)\right].$
+Utilizing Lemma 2, it is straightforward to derive this lemma. The proof is similar to the proof for Lemma 21 in [Cui and Du, 2022b]. ∎
 Theorem 4.
@@
 Appendix B Experiment Details
+Algorithm 4 Pipeline of Preference-Based Multi-agent Reinforcement Learning
+1:Input: Dataset $\mathcal{D}=\{\tau_{i},\tau_{i}^{\prime},\mathbf{y}_{i}\}_{i=1}^{N}$.
+2:Train a agent-wise reward model $r_{\phi}$;
+3:Train a imitation model $\pi_{b}$;
+4:Apply MARL algorithm to learn $\pi_{\mathbf{w}}$ with distribution-based pessimism from $\pi_{b}$;
+5:return $\pi_{\mathbf{w}}$.
 B.1 Implementation Details
+Pipeline
+Our pipeline for Preference-Based Multi-Agent Reinforcement Learning (PbMARL), detailed in Algorithm 4, outlines the key steps for training agents using preference datasets.
+The process begins with training an agent-specific reward model $r_{\phi}$, followed by learning a imitation model $\pi_{b}$.
+The final policy $\pi_{\mathbf{w}}$ is then optimized using a MARL algorithm with distribution-based pessimism derived from $\pi_{b}$.
 Model Configurations
@@
 The MADPO agent network uses RNN to output its Q-values, enabling the agent to make informed decisions based on its observations and learned policies.
+MABCQ and MAIQL, as tested in Section  6, are modified versions of BCQ Fujimoto et al.
+[2019] and IQL Kostrikov et al.
+[2021] tailored for MARL.
+Similar to VDN, the Q-functions in MABCQ and MAIQL are designed to represent the cumulative rewards of all agents collectively.
+This adaptation ensures compatibility with the multi-agent setting while preserving the theoretical and practical foundations of the original algorithms.
+In MAIQL, the coefficient for expectile regression, $\tau$, is consistently set to 0.95 across all experiments.
+For MABCQ, random noise addition is omitted due to the discrete action spaces in the tested environments, and the VAE is replaced with a policy generator trained via imitation learning.
 Table 6 lists the main hyperparameters used in our experiments, while other details can be checked in our codebase: https://github.com/NataliaZhang/marlhf.
 Hyperparameter Default Value
-MSE Loss Coefficient 1
-Reference Log Coefficient 1
+MSE Loss Coefficient $\alpha$ 1 (MPE), 1e-3 (Overcooked)
+Pessimism Coefficient $\beta$ 1 (Spread, Reference), 10 (Tag, Overcooked)
 Prediction Steepness 5
-Number of Environment Timesteps 130
-Number of Environments 128
-Reward Model Type MLP (Spread, Reference), RNN (Tag)
+Episode length 26 (MPE), 400 (Overcooked)
+Reward Model Type MLP (Spread, Reference, Overcooked), RNN (Tag)
 RNN Hidden Size 64 (Tag)
-MLP Layer Dimension 128 (Spread), 64 (Reference)
+MLP Layer Dimension 64
 Reward Model Layers 4
@@
 IL Batch Size 128
-VDN Learning Rate 1e-3
-VDN Epochs 1e4 (Spread, Reference), 1e6 (Tag)
-VDN Batch Size 64
+Policy Model Learning Rate 1e-3
+Policy Model Epochs 1e4 (Spread, Reference), 1e6 (Tag), 1e5 (Overcooked)
+Policy Model Batch Size 128
+MAIQL Expectile Regression Coefficient 0.95
+Optimizer Adam
 Table 6: Main hyperparameters in experiments
@@
 As mentioned in Section 6.2, we collected trajectories in different environments using various policies to ensure a diverse dataset.
-Our dataset size contains 38400 trajectories in each environment.
+Each dataset contains 38400 trajectories in each MPE environment and 960 trajectories in Overcooked.
 The number of trajectory pairs is chosen as 10 times the number of trajectories.
@@
 Note that the original environment is a 1v3 adversary game, and we convert it into a 3-agent cooperative game for better evaluation by fixing the good agent with a MAPPO pretrained policy.
+This environments requires high operation precision.
   • Simple Reference
@@
 The performance of agents is affected particularly under unilateral policy conditions, where misaligned communication signals can significantly impact performance.
+  • Overcooked
+    • Objective: Two agents in a gridworld scoring points by repeatedly completing a three-step process: gathering ingredients, cooking, and serving dishes.
+Each step requires interacting with the environment at specific locations and orientations.
+    • Challenge: The episode length in this environment is notably long (400 timesteps), with sparse reward signals.
+This creates significant challenges for training the reward model, as it may incorrectly attribute rewards to specific behaviors, resulting in inaccurate training objectives and suboptimal learning outcomes.
 These tasks provide a robust framework for evaluating the effectiveness and adaptability of our offline MARLHF algorithm in various multi-agent settings.
@@
 Strengthening regularization (larger $\alpha$) sometimes leads to lower NLL loss, indicating better capability of capturing differences in reward signal.
+Spread-v3 Tag-v3 Reference-v3 Overcooked
+MSE MSE MSE MSE
+Diversified 0.434 1.46 1.19 2.04
+Mix-Unilateral 0.647 1.52 1.09 1.98
+Mix-Expert 0.578 1.78 1.09 2.17
+Pure-Expert 0.673 1.48 2.33 1.72
+Table 9: The mean squared error (MSE) between the standardized predicted rewards and the standardized ground truth rewards.
 In this section, we explore the effects of specific components of our method, focusing on the influence of MSE regularization and the use of diversified datasets.
@@
 As seen in Figure 3, appropriate regularization leads to lower final training loss, suggesting a more stable learning process.
-Furthermore, when analyzing reward prediction accuracy in Figure 4, it becomes evident that the combination of diversified datasets and MSE regularization allows the model to better approximate ground truth rewards, indicating enhanced learning capacity.
+Lower MSE between predicted reward and ground truth (significantly below 2) can indicate a strong correlation between the two (cf.
+9).
+In the MPE experiments, the reward model predictions align closely with the ground truth, and the optimization benefits from dataset diversity are particularly pronounced in the Reference-v3 and Spread-v3 scenarios.
+However, in more complex environments, the reward may exhibit more patterns, making MSE less effective as a metric for assessing reward model quality.
+For example, in the Overcooked environment, assigning rewards for cooking the dish and serving the dish results in very similar returns, as a complete scoring cycle involves both actions, but these two reward function will have squared difference of 2.
 Additionally, our evaluation of the Negative Log-Likelihood (NLL) loss over diversified datasets reveals that stronger regularization can sometimes lead to a reduction in NLL loss, as listed in Table 8.